### v1

# Right-handed neutrino dark matter consistent with the generation of baryon number asymmetry

Abstract
A right-handed neutrino is a promising candidate for dark matter (DM) which has no interaction with nuclei. Since two right-handed neutrinos explain neutrino oscillation data and baryon number asymmetry through seesaw mechanism and leptogenesis based on the out-of-equilibrium decay of the lighter one, the lightest third right-handed neutrino could be free from them and stable to be DM. It suggests a possibility that the right-handed neutrinos could be common key particles for the explanation of three problems in the standard model (SM). We study this possibility by analyzing the structure of a neutrino mass matrix in the scotogenic model for neutrino mass.
1 Introduction
Neutrino oscillation experiments prove the existence of nonzero neutrino mass [1]. Observations of the abundances of light elements and the cosmic microwave background suggest the existence of the baryon number asymmetry in the Universe [2, 3]. However, both the neutrino mass and the baryon number asymmetry cannot be explained in the standard model (SM). It suggests that the SM is incomplete and has to be extended. If the SM is extended by two right-handed neutrinos which are assumed to have Majorana mass, these two problems are considered to be solved through the so called seesaw mechanism [4] and leptogenesis caused by the out-of-equilibrium decay of the right-handed neutrino [5] as long as their Majorana mass is larger than $O(10^{9})$ GeV [6].
On the other hand, the SM has no candidate of dark matter (DM) whose existence has been proved through various astrophysical and cosmological observations. Direct detection experiments of it severely constrain its mass and scattering cross section with a nucleus, and they exclude a lot of DM candidates from a consistency with the required DM relic abundance [7]. It may suggest a possibility that DM has no interaction with nuclei. In this sense, the right-handed neutrino, which is massive and electrically neutral, could be a good candidate of DM since it has no interaction with nuclei. If we combine this idea with the above solution for the neutrino mass and the baryon number asymmetry, the third right-handed neutrino seems to be a good DM candidate if its stability is guaranteed by some symmetry. This stability requirement may impose to introduce an additional doublet scalar $\eta$ since we have to guarantee the existence of their invariant Yukawa couplings between doublet leptons and right-handed neutrinos under the assumed symmetry. As such an example, we have the scotogenic model [8], which can make the neutrino mass generation possible radiatively around TeV scales. The model seems to be very interesting as a simple extension of the SM since three big problems in the SM might be explained by the right-handed neutrinos simultaneously.
However, if we consider such a possibility that the lightest right-handed neutrino $N_{1}$ is DM in the scotogenic model, we have to answer the following hard problems simultaneously:
  1. (i)
    How the required DM relic abundance is realized? If $N_{1}$ has only neutrino Yukawa couplings $h_{\alpha 1}\bar{\ell}_{L_{\alpha}}\eta N_{1}$, large values are required for $h_{\alpha 1}$ to reduce its abundance through the annihilation in the freezeout scenario. This makes it difficult to suppress lepton flavor violating processes (LFV) like $\mu\rightarrow e\gamma$ below the experimental bounds [9].
  2. (ii)
    How the leptogenesis can work successfully in a consistent way with the neutrino mass generation and the DM abundance? If the right-handed neutrinos are supposed to have TeV scale mass, it seems to be difficult to generate sufficient lepton number asymmetry to explain the required baryon number asymmetry [10].
In previous studies aiming to solve these problems including the case where DM is the lightest neutral component of $\eta$, the model is extended by some way, for example, new interaction is introduced to the right-handed neutrinos [11, 12, 13, 14]. In this paper, we propose a solution to these problems within the original scotogenic model without considering any extension of the model.aaaIt is discussed in [15] that $N_{1}$ can be DM consistently with the LFV constraint by considering coannihilation among the right-handed neutrinos without any extension.
The remaining parts are organized as follows. In section 2, we review the model and discuss a neutrino mass matrix derived in it to fix flavor mixing and mass eigenvalues. After that, we study a consistent solution with both the relic DM abundance and the constraint from the lepton flavor violating processes. In section 3, we discuss the consistency of the scenario with successful leptogenesis. We summarize the paper in section 4.
2 Neutrino model with right-handed neutrino DM
A model studied here is a simple extension of the SM with three right-handed neutrinos $N_{k}$ and an inert doublet scalar $\eta$, which is called the scotogenic model [8]. It is assumed to be invariant under an imposed $Z_{2}$ symmetry. New fields $N_{k}$ and $\eta$ are assumed to have its odd parity and its even parity is assigned for all the SM contents. The extended part of Lagrangian is characterized by additional terms
$-{\cal L}_{Y}$ $=$ $\sum_{\alpha,k}h_{\alpha k}\bar{\ell}_{L_{\alpha}}\eta N_{k}+\sum_{k}\frac{1}{2}M_{k}e^{i\gamma_{k}}\bar{N}_{k}N_{k}^{c}+{\rm h.c.},$
$V$ $=$ $\lambda_{1}(\phi^{\dagger}\phi)^{2}+\lambda_{2}(\eta^{\dagger}\eta)^{2}+\lambda_{3}(\phi^{\dagger}\phi)(\eta^{\dagger}\eta)+\lambda_{4}(\phi^{\dagger}\eta)(\eta^{\dagger}\phi)+\frac{\lambda_{5}}{2}\left[(\phi^{\dagger}\eta)^{2}+{\rm h.c.}\right]$ (1)
$+$ $m_{\phi}^{2}\phi^{\dagger}\phi+m_{\eta}^{2}\eta^{\dagger}\eta,$
where $\phi$ is the ordinary Higgs doublet scalar, and $\lambda_{4,5}<0$ and $M_{1}<M_{2}<M_{3}$ are assumed. Neutrino Yukawa couplings $h_{\alpha k}$ are defined using a basis that a mass matrix of right-handed neutrinos is diagonal. Coupling constants in the potential $V$ should satisfy
$\lambda_{1,2}>0,\qquad\lambda_{3}+\lambda_{4}+\lambda_{5}>-\sqrt{\lambda_{1}\lambda_{2}}$ (2)
for the vacuum stability. If $\eta$ is assumed to have no vacuum expectation value (VEV), neutrino mass is not produced at tree level and the $Z_{2}$ symmetry is kept as an exact asymmetry of the model.
After the electroweak symmetry breaking, mass of the components of $\eta$ is fixed through the Higgs VEV as
$M_{\eta^{\pm}}^{2}=m_{\eta}^{2}+\lambda_{3}\langle\phi\rangle^{2},\quad M_{\eta_{R}}^{2}=m_{\eta}^{2}+(\lambda_{3}+\lambda_{4}+\lambda_{5})\langle\phi\rangle^{2},\quad M_{\eta_{I}}^{2}=m_{\eta}^{2}+(\lambda_{3}+\lambda_{4}-\lambda_{5})\langle\phi\rangle^{2},$ (3)
where $\eta^{\pm}$ stands for charged components, and $\eta_{R}$ and $\eta_{I}$ are real and imaginary components, respectively. In this setting, a lightest right-handed neutrino $N_{1}$ or $\eta_{R}$ can be stable due to the $Z_{2}$ symmetry. We suppose that $N_{1}$ is DM and the baryon number asymmetry is generated by leptogenesis through the out-of-equilibrium decay of a heavier right-handed neutrino. In the following study, mass ordering of these extra particles is assumed to satisfy [16]
$M_{1}<M_{\eta_{R}}<M_{2}<M_{3}.$ (4)
2.1 Neutrino mass and flavor mixing
Although neutrino mass is forbidden at tree level in Lagrangian (1), small neutrino mass is induced through a one-loop diagram caused by the Higgs VEV as
$({\cal M}_{\nu})_{\alpha\beta}=\sum_{k}h_{\alpha k}h_{\beta k}\Lambda_{k}e^{i\gamma_{k}}\equiv\sum_{k}\mu_{\alpha\beta}^{k}\Lambda_{k}e^{i\gamma_{k}},$
$\Lambda_{k}=\frac{\lambda_{5}\langle\phi\rangle^{2}}{16\pi^{2}M_{k}}\left[\frac{M_{k}^{2}}{M_{\eta}^{2}-M_{k}^{2}}\left(1+\frac{M_{k}^{2}}{M_{\eta}^{2}-M_{k}^{2}}\ln\frac{M_{k}^{2}}{M_{\eta}^{2}}\right)\right],$ (5)
where $M_{\eta}^{2}=\frac{1}{2}(M_{\eta_{R}}^{2}+M_{\eta_{I}}^{2})$. Since there is no other contribution to the neutrino mass in the model, it is instructive to note an interesting feature of this mass formula.
Its diagonalization matrix can be described by using the Yukawa couplings $h_{\alpha k}$ explicitly. When a condition $\sum_{\alpha}(h_{\alpha j}h_{\alpha k})\propto\delta_{jk}$ is satisfied, which corresponds to the condition that matrices $\mu^{j}$ and $\mu^{k}$ are commutable [11], it is diagonalized by a matrix $U$ as $U^{T}{\cal M}_{\nu}U={\rm diag}(m_{1},m_{2},m_{3})$. This $U$ can be expressed as
$U=(U_{1},U_{2},U_{3}),\quad U_{k}^{T}\equiv(U_{ek},U_{\mu k},U_{\tau k}),\quad U_{\alpha k}\equiv\frac{h_{\alpha k}}{h_{k}},$ (6)
where $h_{k}$ is a real constant. The absolute value of mass eigenvalues $m_{k}$ is fixed as
$m_{k}=\left|\sum_{\alpha}h_{\alpha k}^{2}\right|\Lambda_{k},=p_{k}h_{k}^{2}\Lambda_{k},$ (7)
where $p_{k}$ is defined as $p_{k}=\left|\sum_{\alpha}U_{\alpha k}^{2}\right|$. Here, it should be noted that $\sum_{\alpha}U_{\alpha k}^{2}=1$ is not satisfied if Yukawa couplings $h_{\alpha k}$ are complex. By using this $U$, the PMNS matrix [17] can be written as $V_{\rm PMNS}=U_{e}^{\dagger}UP$ , where $U_{e}$ is a matrix which is used to diagonalize the charged lepton mass matrix and $P$ is a Majorana phase matrix ${\rm diag}(1,e^{i(\gamma_{2}-\gamma_{1})},e^{i(\gamma_{3}-\gamma_{1})})$. Although $V_{\rm PMNS}$ is unitary in a case where $h_{\alpha k}$ is real, it is not unitary in a case where $h_{\alpha k}$ is complex since $U$ is not a unitary matrix.
Since $U$ is the matrix which diagonalizes a symmetric matrix, it may be parametrized as
$U_{\nu}=\left(\begin{array}[]{ccc}c_{12}c_{13}&s_{12}c_{13}&s_{13}\\ -s_{12}c_{23}-c_{12}s_{13}s_{23}&c_{12}c_{23}-s_{12}s_{13}s_{23}&c_{13}s_{23}\\ s_{12}s_{23}-c_{12}s_{13}c_{23}&-c_{12}s_{23}-s_{12}s_{13}c_{23}&c_{13}c_{23}\\ \end{array}\right),$ (8)
where $c_{ij}=\cos\theta_{ij}$ and $s_{ij}=\sin\theta_{ij}e^{i\alpha_{ij}}$. We can represent the neutrino Yukawa couplings $h_{\alpha k}$ and mass eigenvalues $m_{k}$ by using this $U$ as eq. (6) and eq. (7), respectively.
If $h_{\alpha 1}$ has the same structure as $h_{\alpha 2}$ for example, the mass matrix can be diagonalized by $U$ given in eq. (8) but we have a zero mass eigenvalue $m_{1}=0$. On the $m_{2}$ formula in eq. (7), $p_{2}h_{2}^{2}\Lambda_{2}$ is replaced by $p_{1}h_{1}^{2}\Lambda_{1}+p_{2}h_{2}^{2}\Lambda_{2}$. As such a concrete example, neutrino Yukawa couplings may be taken as [14]
$h_{ei}=h_{\mu i}=-h_{\tau i}=\frac{1}{\sqrt{3}}h_{i}~{}~{}(i=1,2),\quad h_{e3}=0,~{}h_{\mu 3}=h_{\tau 3}=\frac{h_{3}}{\sqrt{2}}.$ (9)
The mixing matrix $U$ can be taken as the tribimaximal mixing matrix [18] by fixing the mixing angles as
$s_{12}=\frac{1}{\sqrt{3}},\quad s_{23}=\frac{1}{\sqrt{2}},\quad s_{13}=0,\quad\alpha_{12}=\alpha_{23}=0,$ (10)
and the neutrino mass eigenvalues are obtained as
$m_{1}=0,\quad m_{2}=h_{1}^{2}\Lambda_{1}+h_{2}^{2}\Lambda_{2},\quad m_{3}=h_{3}^{2}\Lambda_{3}.$ (11)
Since the fault of the tribimaximal mixing matrix as the PMNS matrix could be overcome by considering a nontrivial flavor mixing matrix $U_{e}$ in the charged lepton sector [14], it can become a phenomenologically interesting possibility.
Neutrino oscillation data fix squared mass differences $\Delta m_{32}^{2}\equiv m_{3}^{2}-m_{2}^{2}$ and $\Delta m_{21}^{2}\equiv m_{2}^{2}-m_{1}^{2}$. If $h_{1}$ is sufficiently small and then $m_{1}\simeq 0$, Yukawa couplings $h_{2}$ and $h_{3}$ can be determined by applying eq. (7) to
$m_{2}=\sqrt{\Delta m_{21}^{2}},\qquad m_{3}=\sqrt{\Delta m_{32}^{2}+\Delta m_{21}^{2}}.$ (12)
If we use the best fit values from neutrino oscillation data given in [19] for $\sin\theta_{ij}$ in the matrix $U$, we can calculate two effective masses relevant to two observables, that is, the effective masses for the neutrinoless double $\beta$ decay and the $\beta$ decay. They are predicted to be
$m_{\beta\beta}\simeq\left|\sum_{j=1}^{3}U_{ej}^{2}m_{j}\right|=0.0037~{}{\rm eV},\quad m_{\beta}\simeq\sqrt{\sum_{j=1}^{3}|U_{ej}|^{2}m_{j}^{2}}=0.0088~{}{\rm eV}.$ (13)
These results are much smaller than their present bounds [20, 21].
[…]
$M_{1}=1.5~{}{\rm TeV},\quad M_{\eta_{R}}\simeq M_{2}\simeq M_{3}\simeq 3~{}{\rm TeV},\quad\lambda_{5}=-10^{-3},$
$h_{1}=6\times 10^{-9},\quad h_{2}=1.11\times 10^{-3},\quad h_{3}=4.15\times 10^{-3},$ (14)
where we assume $\alpha_{12}=\alpha_{13}=0,~{}\alpha_{23}=\frac{\pi}{2}$ and use the best fit values from neutrino oscillation data for $\sin\theta_{ij}$. A smaller $|\lambda_{5}|$ makes $h_{k}$ larger for fixed values of $M_{k}$. Since DM is supposed to be $N_{1}$ here, there is no constraint on a lower bound of the value of $|\lambda_{5}|$ differently from the case where $\eta_{R}$ is DM [10]. We find that the present bounds of LFV like $\mu\rightarrow e\gamma$ [19] can be satisfied for these parameters. We come back to this point later. In the DM study, it may be useful to note that the above values of $h_{2,3}$ are not affected even if $h_{1}$ takes the same order value as $h_{2,3}$ as long as $M_{1}$ is much smaller than $M_{\eta_{R}}$. It is caused by the fact that $\Lambda_{1}$ has a suppression of $O(M_{1}/M_{\eta_{R}})$ compared with $\Lambda_{2,3}$. It allows $N_{1}$ with keV scale mass to have an acceptable relic abundance as DM for the same $h_{2,3}$ given in eq. (14).
Finally, we note an important consequence caused by the feature of this mass matrix addressed above. If we consider leptogenesis based on the out-of-equilibrium decay of the right-handed neutrino $N_{k}$, the model predicts an interesting result. $CP$ asymmetry in this process is known to be caused by the interference between tree and one-loop diagrams, and it should be nonzero at least to make leptogenesis work. Since it is proportional to ${\rm Im}(\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j})^{2}$ for $k\not=j$ as shown in eqs. (21) and (25) later, this condition cannot be satisfied because of $\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j}\propto\delta_{kj}$ for real neutrino Yukawa couplings $h_{\alpha k}$ even if Majorana phases in the right-handed neutrino mass are taken into account. Thus, Yukawa couplings $h_{\alpha k}$ should be complex, and the PMNS matrix cannot be unitary in that case. However, we should note that there is an exceptional case where the $CP$ asymmetry can take a nonzero value even if Yukawa couplings $h_{\alpha k}$ are real. The existence of a zero mass eigenvalue presents such a case as found in an example given in eq. (9) when Majorana phases exist in the right-handed neutrino mass [14]. The PMNS matrix becomes unitary in this case.
2.2 Relic abundance of right-handed neutrino DM
Right-handed neutrino $N_{1}$ is stable to be DM under the condition (4). However, its abundance is difficult to reach the required value in the freezeout scenario as mentioned in the introduction. It can reach the equilibrium at the temperature higher than its mass through both the $\eta$ decay and the scatterings caused by the neutrino Yukawa couplings $h_{\alpha 1}$ if they have sufficient magnitude. In that case, $h_{\alpha 1}$ has to take a large value of $O(1)$ to reduce its abundance sufficiently through the annihilation. Unfortunately, such a large Yukawa coupling causes LFV processes like $\mu\rightarrow e\gamma$ at an unacceptable level through the one-loop diagram [9]. Since this difficulty is caused by a feature that $N_{1}$ has interaction due to the neutrino Yukawa couplings only, new interaction like $U(1)^{\prime}$ has been introduced to reduce its abundance in many articles [11, 12].
In the present model, since the neutrino Yukawa coupling $h_{1}$ is supposed to be very small as shown in eq. (14), its thermal abundance is originally difficult to reach a value larger than the required one so that the freezeout scenario cannot be applied to the model. However, even in that case, we can consider an alternative possibility such that it is realized through the decay of $\eta_{R}$ following the freezein scenario [22]. $N_{1}$ is produced through the $\eta_{R}$ decay caused by the neutrino Yukawa coupling $h_{1}$. Since $h_{1}$ is supposed to be sufficiently small here, the $N_{1}$ generation through the $\eta_{R}$ decay is expected to be largely suppressed. On the other hand, the next lightest DM candidate $\eta_{R}$ is in the thermal equilibrium through the SM interactions. Decrease of the number density of $\eta_{R}$ is controlled by both the coannihilation [23] to the SM contents and the decay to $N_{1}$. The $N_{1}$ relic abundance is determined by both the $N_{1}$ freezein and the $\eta_{R}$ freezeout depending on their relative effectiveness.
We express the number density of $\eta_{R}$ and $N_{1}$ in the comoving volume by $Y_{i}=\frac{n_{i}}{s}$, where $n_{i}$ is the number density of particle species $i$ and $s$ is the entropy density. If we define a dimensionless parameter $z_{2}$ as $z_{2}\equiv\frac{M_{2}}{T}$ and represent an equilibrium value of $Y_{i}$ as $Y_{i}^{\rm eq}$, Boltzmann equations for $Y_{\eta_{R}}$ and $Y_{N_{1}}$ can be written as
$\frac{dY_{\eta_{R}}}{dz_{2}}=-\frac{z_{2}}{H(M_{2})s}\left\{2\gamma^{S}(\eta_{R}\eta_{R})+\gamma^{D}(\eta_{R})\right\}\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right),$ (15)
$\frac{dY_{N_{1}}}{dz_{2}}=\frac{z_{2}}{H(M_{2})s}\gamma^{D}(\eta_{R})\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right),$ (16)
where $H(M_{2})$ represents the Hubble parameter at $T=M_{2}$. Reaction densities $\gamma^{D}(a)$ and $\gamma^{S}(ab)$ are respectively related to thermally averaged decay width $\langle\Gamma_{a}\rangle$ for $a\rightarrow ij$ and cross section $\langle\sigma_{ab}|v|\rangle$ for scattering $ab\rightarrow ij$ as[25]
$\gamma^{D}(a)=\langle\Gamma_{a}\rangle n_{a}^{\rm eq},\qquad\gamma^{S}(ab)=\langle\sigma_{ab}|v|\rangle n_{a}^{\rm eq}n_{b}^{\rm eq},$ (17)
where $n_{a}^{\rm eq}$ is the equilibrium number density of a particle $a$. The present DM abundance can be expressed as $\Omega_{DM}=M_{DM}Y_{DM}^{\infty}s_{0}/\rho_{0}$, where $\rho_{0}$ and $s_{0}$ are the energy density and the entropy density in the present Universe. They are given as $\rho_{0}=3M_{\rm pl}^{2}H_{0}^{2}$ and $s_{0}=2.27\times 10^{-38}$ GeV3, and then it determines the required $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ through $\Omega_{N_{1}}h^{2}=0.12$ as $Y_{N_{1}}^{\infty}=4.2\times 10^{-10}(M_{1}/{\rm GeV})^{-1}$.
Here, we roughly estimate both the freezeout temperature $T_{\eta}$ of the coannihilation of $\eta_{R}$ and a value of $h_{1}$ for the $\eta_{R}$ decay to reach the equilibrium at $T_{\eta}$. The freezeout temperature $T_{\eta}(\equiv M_{\eta_{R}}/z_{\eta})$ can be estimated through a condition $2n_{\eta_{R}}^{\rm eq}(T_{\eta})\langle\sigma_{\eta\eta}|v|\rangle=H(T_{\eta})z_{\eta}$ by assuming instantaneous freezeout as found from eq. (15). Using the Hubble parameter $H(T)$ given in eq. (19), $z_{\eta}$ can be determined by solving this condition as [26]
$z_{\eta}=\ln\frac{0.384M_{\rm pl}\langle\sigma_{\eta\eta}|v|\rangle_{\eta}M_{\eta_{R}}}{(g_{\ast}z_{\eta})^{1/2}}.$ (18)
The $h_{1}$ for the $\eta_{R}$ decay to reach the equilibrium at $T_{\eta}$ can be estimated from the condition $\Gamma_{\eta_{R}}=H(T_{\eta})$. The $\eta_{R}$ decay width $\Gamma_{\eta_{R}}$ and the Hubble parameter $H(T)$ are given as
$\Gamma_{\eta_{R}}=\frac{q_{1}h_{1}^{2}}{32\pi}M_{\eta_{R}}\left(1-\frac{M_{1}^{2}}{M_{\eta_{R}}^{2}}\right)^{2},\quad H(T)=\left(\frac{\pi^{2}}{90}g_{\ast}\right)^{1/2}\frac{T^{2}}{M_{\rm pl}},$ (19)
where $M_{\rm pl}$ is the reduced Planck mass and $q_{k}$ is defined as $q_{k}=\sum_{\alpha}(U_{\alpha k}^{\ast}U_{\alpha k})$. This condition determines $h_{1}$ as
$h_{1}c_{p}=3\times 10^{-8}\left(\frac{M_{\eta_{R}}}{3~{}{\rm TeV}}\right)^{1/2},$ (20)
where $c_{p}$ is a constant determined by the phase space volume.cccWhen the mass of $\eta_{R}$ and $N_{1}$ take closely degenerate values, $c_{p}$ can be represented as $c_{p}=2\delta_{1}$, where $\delta_{1}$ is defined by $M_{1}=(1-\delta_{1})M_{\eta_{R}}$ with $\delta_{1}\ll 1$.
If $h_{1}$ takes a larger value than the one given in eq. (20), the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be approximated well as $Y_{N_{1}}^{\infty}=Y_{\eta_{R}}^{\rm eq}(T_{c})$ through the freezein caused by the $\eta_{R}$ decay, where $T_{c}$ is defined by $Y_{N_{1}}(T_{c})=Y_{\eta_{R}}^{\rm eq}(T_{c})$ since the $\eta_{R}$ decay at $T<T_{c}$ does not affect $Y_{N_{1}}^{\infty}$ substantially. $Y_{N_{1}}^{\infty}$ is determined by the coupling $h_{1}$ for a fixed $M_{1}$ and the freezeout of the $\eta_{R}$ (co)annihilation does not affect the result. However, this situation could change when $h_{1}$ takes a smaller value than the one given in eq. (20) and $T_{c}<T_{\eta}$ is satisfied. In this case, the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be affected by the equilibrium number density of $\eta_{R}$ at $T_{\eta}$. $T_{\eta}$ has been known to be determined through eq. (18) by the couplings $\lambda_{3}$ and $\lambda_{4}$ through the study of the inert doublet DM model [24]. If we use this knowledge, the required $N_{1}$ relic abundance could be predicted through a relation $Y_{N_{1}}^{\infty}\simeq Y_{\eta_{R}}^{\rm eq}(T_{\eta})$. It also gives constraints on the couplings $\lambda_{3}$ and $\lambda_{4}$.
Fig. 1   Left panel: Evolution of the number density of $N_{1}$ and $\eta_{R}$ in the comoving volume for three values of $M_{1}$ are plotted with red lines, 1.5 TeV (solid), 150 GeV (dashed) and 1.5 GeV (dotted). $h_{1}$ is fixed at $7.8\times 10^{-9}$, $4.5\times 10^{-8}$ and $7.0\times 10^{-6}$ for each $M_{1}$, respectively. $M_{\eta_{R}}=3$ TeV and $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$ are assumed. Horizontal black lines represent $Y_{N_{1}}$ values required for each value of $M_{1}$ to realize the present DM relic density. While an $Y_{\eta_{R}}$ line for $M_{1}=1.5$ TeV deviates from $Y_{\eta_{R}}^{\rm eq}$, $Y_{\eta_{R}}$ in other two cases coincides with $Y_{\eta_{R}}^{\rm eq}$ since the $\eta_{R}$ decay is already in the equilibrium at a temperature higher than $T_{c}$. Right panel: Contours of $Y_{\eta_{R}}(T_{\eta})$ for $M_{\eta_{R}}=3$ TeV in the $(\lambda_{3},\lambda_{4})$ plane. Each red line represents $Y_{\eta_{R}}(T_{\eta})=10^{-10.5},10^{-11},10^{-11.5},10^{-11.75}$, and $10^{-12}$ outward. A dashed line represents $\lambda_{3}+\lambda_{4}+\lambda_{5}=-\sqrt{\lambda_{1}\lambda_{2}}$ in the case of $\lambda_{2}=1.0$, and an upper region of it satisfies the potential stability. No constraint comes from direct detection experiments on $\lambda_{3,4}$ since DM is $N_{1}$.
In order to show this observation concretely, we solve the Boltzmann equations for $Y_{N_{1}}$ and $Y_{\eta_{R}}$ using the parameters given in eq. (14) and $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$. Since $M_{1}$ and $h_{1}$ can be free from neutrino mass determination as addressed below eq. (14), we plot the solutions in the left panel for the cases $(M_{1},h_{1})=(1.5~{}{\rm TeV},7.8\times 10^{-9})$, $(150~{}{\rm GeV},4.5\times 10^{-8})$, and $(1.5~{}{\rm GeV},7.0\times 10^{-6})$ where $h_{1}$ is tuned to give the required DM abundance. In this panel, $Y_{N_{1}}$ required to explain the DM abundance is shown by horizontal black lines for each case. In the latter two cases, the required $Y_{N_{1}}^{\infty}$ is realized through the pure freezein due to the $\eta_{R}$ decay. $Y_{\eta_{R}}$ does not deviate from the equilibrium one since its decay to $N_{1}$ is already in the equilibrium at $T>T_{\eta}$ and then $Y_{\eta_{R}}=Y_{\eta_{R}}^{\rm eq}$ is kept. The freezeout of the $\eta_{R}$ coannihilation in a case of $M_{\eta_{R}}=3$ TeV is shown in the right panel, in which contours for $Y_{\eta_{R}}^{\rm eq}(T_{\eta})=10^{-10.5},10^{-11},10^{-11.5},10^{-11.75}$, and $10^{-12}$ are plotted in the $(\lambda_{3},\lambda_{4})$ plane. Since $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$ is assumed in the left panel, we find $Y_{\eta_{R}}^{\rm eq}(T_{\eta})=10^{-11.75}$ is relevant. It is smaller than the required $Y_{N_{1}}^{\infty}$ in these cases, and then the freezeout of the $\eta_{R}$ coannihilation is irrelevant to the estimation of $Y_{N_{1}}^{\infty}$. As an extreme case of these examples, we may consider $N_{1}$ with keV scale mass which can be consistent with the neutrino oscillation data as discussed in the previous part. We can confirm that the required relic DM abundance is obtained by tuning $h_{1}$ suitably, for example, $h_{1}=3.7\times 10^{-3}$ for $M_{1}=5$ keV.
On the other hand, the freezeout of the $\eta_{R}$ coannihilation is relevant to the estimation of $Y_{N_{1}}$ in the case of $M_{1}=1.5$ TeV. Although $Y_{N_{1}}(T_{c})$ is smaller than the required value, the late time decay of the frozen $\eta_{R}$ contributes to $Y_{N_{1}}$ and make it the required value $Y_{N_{1}}^{\infty}$ as found from the figure. Since the stability of the potential $V$ in eq. (1) requires $\lambda_{3,4}$ should satisfy the condition (2), an allowed region of $(\lambda_{3},\lambda_{4})$ is a part over the dashed line in the right panel. Although $Y_{N_{1}}$ is determined by the values of $\lambda_{3}$ and $\lambda_{4}$ in the $M_{1}=1.5$ TeV case, we note that there is no constraint on them from direct detection experiments since $N_{1}$ is DM which cannot interact with nuclei. It is a crucial difference from the case where $\eta_{R}$ is DM. The coupling constants $\lambda_{3}$ and $\lambda_{4}$ could be also constrained through the correction to the Higgs mass caused by them. This constraint is reported to be satisfied for $M_{\eta_{R}}\leq 4.4$ TeV in [27].
In the above study, we do not take into account the effect caused by the simultaneous explanation of the baryon number asymmetry in the Universe through leptogenesis based on the $N_{2}$ decay. Since it could produce $\eta_{R}$ at late time, the $N_{1}$ abundance could be affected by it if the $N_{2}$ decay occurs near the freezeout of $\eta_{R}$. We discuss this point in the next part.
3 Consistency with leptogenesis
In the present model, the baryon number asymmetry is considered to be produced through leptogenesis caused by the out-of-equilibrium decay of $N_{2}$. This requires that $N_{2}$ should be in the thermal bath and its decay to $\eta_{R}$ occurs in a $CP$ asymmetric way at a temperature higher than 100 GeV where sphaleron interaction is in the equilibrium. Values of $h_{2,3}$ given in eq. (14) constrained by the neutrino oscillation data are large enough to produce $N_{2,3}$ in the thermal bath as relativistic particles through the inverse decay and scatterings as long as reheating temperature is larger than $M_{3}$. $CP$ asymmetry in the decay $N_{2}\rightarrow\ell_{\alpha}\eta$ can be caused through an interference between a tree diagram and a one-loop diagram mediated by $N_{1}$ and $N_{3}$.
The $CP$ asymmetry caused by one loop vertex diagrams is estimated through the formula
$\varepsilon=\frac{1}{8\pi}\sum_{k=1,3}\frac{{\rm Im}\left[\left(\sum_{\alpha}h_{\alpha 2}h^{\ast}_{\alpha k}\right)^{2}\right]}{\sum_{\alpha}h_{\alpha 2}h_{\alpha 2}^{\ast}}F\left(\frac{M_{k}^{2}}{M_{2}^{2}},\frac{M_{\eta_{R}}^{2}}{M_{2}^{2}}\right)$
$F(x,y)=\sqrt{x}\left[1-\frac{x-2y+1}{(1-y)^{2}}\ln\left(\frac{x-2y+1}{x-y^{2}}\right)\right],$ (21)
where $M_{2}\simeq M_{\eta_{R}}$ is taken into account. If we note that $h_{1}$ is supposed to be very small for the DM explanation as discussed in the previous part, the dominant contribution to this $\varepsilon$ is found to be caused by $N_{3}$ and it can be estimated by using eqs. (7) and (12) asdddIt should be noted that the existence of both a zero mass eigenvalue and Majorana phases of the right-handed neutrino masses are required to guarantee a non-zero $\varepsilon$ when neutrino Yukawa couplings $h_{\alpha k}$ are real [14].
$\varepsilon$ $\simeq$ $\frac{h_{3}^{2}}{8\pi}\frac{{\rm Im}[(\sum_{\alpha}U_{\alpha 2}U_{\alpha 3}^{\ast})^{2}e^{i(\gamma_{2}-\gamma_{3})}]}{\sum_{\alpha}U_{\alpha 2}U_{\alpha 2}^{\ast}}F\left(\frac{M_{3}^{2}}{M_{2}^{2}},\frac{M_{\eta_{R}}^{2}}{M_{2}^{2}}\right)$ (22)
$\simeq$ $4.7\times 10^{-7}F\left(\frac{M_{3}^{2}}{M_{2}^{2}},\frac{M_{\eta_{R}}^{2}}{M_{2}^{2}}\right),$
where parameters in eq. (14) and $\sin\theta_{ij}$ and $\alpha_{ij}$ written below it are used, and $\sin(\gamma_{2}-\gamma_{3})=1$ is assumed in the second line. We should note that a smaller $|\lambda_{5}|$ can make Yukawa couplings larger and then $\varepsilon$ larger but it makes the out-of-equilibrium decay of $N_{2}$ harder.
The $N_{2}$ decay is in the equilibrium at temperature $T\geq T_{D}$. $T_{D}$ is defined by $\Gamma_{N_{2}}=H(T_{D})$ where the decay width of $N_{2}$ can be expressed as
$\Gamma_{N_{2}}=\frac{q_{2}h^{2}_{2}}{16\pi}M_{2}\left(1-\frac{M_{\eta_{R}}^{2}}{M_{2}^{2}}\right)^{2}.$ (23)
If the washout processes of the generated lepton number asymmetry are frozen out at temperature $T_{F}$, the lepton number asymmetry caused by the out-of-equilibrium decay of $N_{2}$ could escape the washout in the case $T_{D}<T_{F}$ since the dominant part of lepton number asymmetry is expected to be generated at $T~{}{{}^{<}_{\sim}}~{}T_{D}$. Such situation could happen if the $N_{2}$ decay is forced to be delayed and be kept in the out-of-equilibrium by making $T_{D}$ to a lower temperature. On the other hand, the out-of-equilibrium decay of $N_{2}$ is favored to start at the temperature $T~{}{{}^{>}_{\sim}}~{}M_{2}$ for the efficient generation of the lepton number asymmetry since the $N_{2}$ abundance can keep a large value as the relativistic particle there.
Although $h_{2}$ fixed by the neutrino oscillation data cannot realize such a situation for a TeV scale $M_{2}$, it could be realized through an additional suppression factor for $\Gamma_{N_{2}}$. It is caused the phase space suppression due to the strict degeneracy between $M_{2}$ and $M_{\eta_{R}}$ [16]. In fact, if their degeneracy is expressed as $M_{\eta_{R}}=(1-\delta_{2})M_{2}$ with $\delta_{2}\ll 1$, $\Gamma_{N_{2}}$ is found to be rewritten with an additional suppression factor $\delta_{2}$ as
$\Gamma_{N_{2}}\simeq\frac{h_{2}^{2}\delta_{2}^{2}}{4\pi}M_{2}=2.9\times 10^{-12}\left(\frac{M_{2}}{3~{}{\rm TeV}}\right)\left(\frac{\delta_{2}}{10^{-4}}\right)^{2}~{}{\rm GeV},$ (24)
where we use the values for $\sin\theta_{ij}$ and $\alpha_{ij}$ given below eq. (14). Since it shows that $\Gamma_{N_{2}}<H(M_{2})$ is satisfied for $\delta_{2}<10^{-4}$ at $M_{2}=3$ TeV, $N_{2}$ is expected to be out-of-equilibrium at $T~{}{{}^{>}_{\sim}}~{}M_{2}$. Thus, if the condition $T_{D}<T_{F}$ is fulfilled, the generated lepton number asymmetry $Y_{L}$ defined as $Y_{L}\equiv Y_{\ell}-Y_{\bar{\ell}}$ could be estimated as $Y_{L}=\varepsilon Y_{N_{2}}^{\rm eq}$ where $Y_{N_{2}}^{\rm eq}$ stands for the equilibrium value of $N_{2}$ as the relativistic particle. In that case, lepton number asymmetry $Y_{L}=O(10^{-10})$ required for the explanation of baryon number asymmetry in the Universe is expected to be obtained if $\varepsilon$ takes a value of $O(10^{-7})$ in eq. (22). Unfortunately, since $\delta_{2}\ll 1$ also suppresses $\varepsilon$ heavily as found from eq.(21), this scenario cannot work.
If $M_{2}\simeq M_{1}$ or $M_{3}$ is satisfied, the $CP$ asymmetry $\varepsilon$ could be enhanced by the resonance due to a self-energy diagram mediated by $N_{1}$ or $N_{3}$. It is expressed as [28, 29, 30]
$\varepsilon$ $=$ $\frac{{\rm Im}\left[\left(\sum_{\alpha}h_{\alpha 2}h^{\ast}_{\alpha k}\right)^{2}\right]}{(\sum_{\alpha}h_{\alpha 2}h_{\alpha 2}^{\ast})(\sum_{\alpha}h_{\alpha k}h_{\alpha k}^{\ast})}\frac{(M_{2}^{2}-M_{k}^{2})M_{2}\Gamma_{N_{k}}}{(M_{2}^{2}-M_{k}^{2})^{2}+M_{2}^{2}\Gamma_{N_{k}}^{2}},$ (25)
where $k=1,3$ and $\Gamma_{N_{k}}$ is the decay width of $N_{k}$. Since $N_{1}$ is stable in the present model and then $\Gamma_{N_{1}}=0$, only $N_{3}$ can contribute to the resonant $CP$ asymmetry. Thus, $\varepsilon$ can be written by using the values for $\sin\theta_{ij}$ and $\alpha_{ij}$ given below eq. (14) as
$\varepsilon=-{\rm Im}\left[\left(\sum_{\alpha}U_{\alpha 2}U^{\ast}_{\alpha 3}\right)^{2}e^{i(\gamma_{2}-\gamma_{3})}\right]\frac{2\delta_{3}\tilde{\Gamma}_{N_{3}}}{4\delta_{3}^{2}+\tilde{\Gamma}_{N_{3}}^{2}}=0.67\frac{2\delta_{3}\tilde{\Gamma}_{N_{3}}}{4\delta_{3}^{2}+\tilde{\Gamma}_{N_{3}}^{2}},$ (26)
where $M_{3}=(1+\delta_{3})M_{2}$ and $\tilde{\Gamma}_{N_{3}}=\Gamma_{N_{3}}/M_{2}$, and $\sin(\gamma_{2}-\gamma_{3})=1$ is assumed. If $\delta_{2}>\delta_{3}$ is satisfied, $\tilde{\Gamma}_{N_{3}}$ can be estimated as $\tilde{\Gamma}_{N_{3}}=h_{3}^{2}\delta_{2}^{2}/4\pi$ and then we have $\varepsilon\simeq 0.027h_{3}^{2}\delta_{2}^{2}/\delta_{3}$ for $\tilde{\Gamma}_{N_{3}}<2\delta_{3}$.
It is crucial to know what value of $\delta_{2}$ can make the condition $100~{}{\rm GeV}<T_{D}<T_{F}$ be satisfied for the parameters given in eq. (14). By calculating the relevant reaction rates of washout processes, we find that $T_{F}~{}{{}^{>}_{\sim}}~{}0.1M_{2}$ so that $\delta_{2}<3\times 10^{-5}$ guarantees $T_{D}<T_{F}$. This suggests that the $CP$ asymmetry $\varepsilon$ can take a value of $O(10^{-7})$ in the case $\delta_{3}=O(10^{-9})$, which makes successful leptogenesis possible as addressed below eq. (24). However, if the number density of $N_{2}$ does not decrease to the required $Y_{N_{1}}^{\infty}$, the $\eta_{R}$ produced through the $N_{2}$ decay contributes to the $N_{1}$ relic density. Thus, the estimation of the $N_{1}$ relic abundance in the previous section cannot naively applied to this scheme for leptogenesis. The effect of $\eta_{R}$ produced in the late time $N_{2}$ decay needs to be taken into account for the estimation of the $N_{1}$ relic abundance.
In order to examine its effect to the $N_{1}$ relic abundance, we solve numerically Boltzmann equations for $Y_{N_{2}}$, $Y_{L}$, $Y_{\eta_{R}}$ and $Y_{N_{1}}$. Using a dimensionless parameter $z_{2}$, they are given as
$\frac{dY_{N_{2}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left\{\gamma^{S}(N_{2}\ell)+\gamma^{S}(N_{2}\eta)+\gamma^{D}(N_{2})\right\}\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right),$
$\frac{dY_{L}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\varepsilon\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)+\left\{\sum_{k=2,3}\gamma^{D}(N_{k})+2(\gamma^{S}(\ell\eta)+\gamma^{S}(\ell\ell))\right\}\frac{Y_{L}}{Y_{\ell}^{\rm eq}}\right],$
$\frac{dY_{\eta_{R}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\left\{2\gamma^{S}(\eta_{R}\eta_{R})+\gamma^{D}(\eta_{R})\right\}\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right)+\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)\right],$
$\frac{dY_{N_{1}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})}\gamma^{D}(\eta_{R})\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right),$ (27)
where $\gamma^{S}({N_{2}\ell})$ and $\gamma^{S}(N_{2}\eta)$ are reaction densities of the scattering $N_{3}\ell\leftrightarrow N_{2}\ell$ and $N_{3}\eta\leftrightarrow N_{2}\eta$ for $N_{2}$ production, and $\gamma^{S}(\ell\eta)$ and $\gamma^{S}(\ell\ell)$ are the ones for the lepton number violating scattering $\ell\eta\leftrightarrow\bar{\ell}\eta^{\dagger}$ and $\ell\ell\leftrightarrow\eta\eta$. A new term is introduced to the equation for $Y_{\eta_{R}}$ to take account of the contribution from the late time decay of $N_{2}$ in comparison with the original one in eq. (15). Change of the $Y_{N_{1}}$ evolution caused by imposing the production of the required lepton number asymmetry can be found from the solutions.
Their solutions for the same values of $M_{1}$ and $(\lambda_{3},\lambda_{4})$ used in Fig. 1 are plotted as functions of $z_{2}$ in the left panel of Fig. 2. We use the values of $M_{\eta_{R}}$, $M_{2}$ and $M_{3}$ given in (14) with $\delta_{2}=3\times 10^{-5}$ and $\delta_{3}=10^{-9}$ which are adopted referring to the above discussion. The figure shows that the lepton number asymmetry $|Y_{L}|$ plotted by an orange line is generated to converge to a final value at $z_{2}\sim 10$, which is larger compared with a required range to explain the baryon number asymmetry shown by dotted yellow lines. Since the resulting lepton number asymmetry takes a value $Y_{L}\simeq\varepsilon Y_{N_{2}}^{\rm eq}$ where $Y_{N_{2}}^{\rm eq}$ is the one of the relativistic particles, the washout of the generated lepton number asymmetry is found to be sufficiently suppressed as a result of the late time decay of $N_{2}$ as expected from the above discussion.
Fig. 2  Evolution of the lepton number asymmetry $|Y_{L}|$ and the $N_{1}$ number density $Y_{N_{1}}$ in the comoving volume. $|Y_{L}|$ is plotted by a orange line, which is caused by the out-of-equilibrium decay of $N_{2}$. $Y_{N_{2}}^{\rm eq}$ should be noted to coincide with $Y_{\eta_{R}}^{\rm eq}$ since $M_{2}\simeq M_{\eta_{R}}$ is assumed. Yellow dashed horizontal lines represent a region of $|Y_{L}|$ to explain the required baryon number asymmetry. In the left panel, $Y_{N_{1}}$ is plotted by red lines for the same $M_{1}$ as the ones in the left panel of Fig. 1. Only $h_{1}$ is tuned to realized the present DM abundance which is represented by black solid, dashed and dotted lines for each $M_{1}$. Other parameters are fixed to the same values as the ones in Fig. 1. In the right panel, $(\lambda_{3},\lambda_{4})$ is changed to $(2.5,-1.0)$ from the ones used in the left panel keeping the $h_{1}$ value unchanged. While $Y_{N_{1}}$ becomes a little bit smaller than the required value $Y_{N_{1}}^{\infty}$ in $M_{1}=150$ GeV and 1.5 GeV, $Y_{N_{1}}+Y_{\eta_{R}}=Y_{N_{1}}^{\infty}$ is realized for $M_{1}=1.5$ TeV at $z_{2}=300$. It means that $Y_{N_{1}}=Y_{N_{1}}^{\infty}$ is obtained finally through the $\eta_{R}$ decay.
On the other hand, we find from the figure that $Y_{\eta_{R}}$ leaves $Y_{\eta_{R}}^{\rm eq}$ around $z_{2}=20$, which is considered to be caused due to the break of the balance between the $\eta_{R}$ production by the $N_{2}$ decay and its disappearance due to both the decay and the coannihilation. It shows that the late time additional $\eta_{R}$ production through the $N_{2}$ decay affects the estimation of the $N_{1}$ relic abundance. It forces us to change the strength of the $\eta_{R}$ decay and the coannihilation of $\eta_{R}$ in order to realize the required $N_{1}$ abundance since we impose simultaneous explanation of the DM abundance and the baryon number asymmetry on the model.eeeSince the $N_{2}$ decay is relevant to the leptogenesis, we do not change the parameters relevant to it here. In the figure, $Y_{\eta_{R}}$ and $Y_{N_{1}}$ are plotted by using blue and red lines, respectively. Different line types are used for different masses, $M_{1}$=1.5 TeV (solid), 150 GeV (dashed), and 1.5 GeV (dotted). $Y_{N_{1}}^{\infty}$ which can reproduce $\Omega_{N_{1}}h^{2}=0.12$ as the present DM abundance is shown by horizontal black solid, dashed and dotted lines for each $M_{1}$. Since $h_{1}$ does not affect the values of $h_{2,3}$ determined by the neutrino oscillation data as long as $h_{1}$ is small enough, it can be tuned to result in the required $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$. By tuning $h_{1}$ suitably to $h_{1}=2.3\times 10^{-9}$ and $2.3\times 10^{-8}$ for each case $M_{1}=$150 GeV and 1.5 GeV, $Y_{N_{1}}$ is found to converge to the required $Y_{N_{1}}^{\infty}$ at $z_{2}$ where $Y_{N_{1}}=Y_{N_{2}}$ is realized. These values of $h_{1}$ are much smaller than the ones used in Fig. 1. It is considered to be caused to cancel the effect of $\eta_{R}$ additionally produced through the $N_{2}$ decay by making the relative strength of the coannihilation of $\eta_{R}$ to the $\eta_{R}$ decay larger. The similar feature can be observed in the evolution of the number density of the keV scale $N_{1}$ also. For example, the DM relic abundance requires $h_{1}=6.5\times 10^{-5}$ for $M_{1}=5$ keV, and these parameters are confirmed not to affect the result of leptogenesis substantially. On the other hand, we find that the relation $Y_{N_{1}}+Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ cannot be changed by tuning the value of $h_{1}$ in the case $M_{1}=1.5$ TeV. This may be understood if we note that $Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ at $z_{2}=300$ where the coannihilation of $\eta_{R}$ is considered to be frozen out. It suggests that the freezeout of the coannihilation of $\eta_{R}$ plays an important role and then $\lambda_{3}$ and $\lambda_{4}$ have to be tuned especially when $M_{1}$ takes a similar order value to $M_{\eta_{R}}$.
In the right panel, $Y_{N_{1}}$ and $Y_{\eta_{R}}$ are plotted for the case that $\lambda_{3}$ and $\lambda_{4}$ are shifted from the ones in the left panel to $(\lambda_{3},\lambda_{4})=(2.5,-1.0)$ keeping other parameters unchanged. While $Y_{N_{1}}$ in the case of $M_{1}=150$ GeV and 1.5 GeV is found to take a little bit smaller value than the one in the left panel, $Y_{N_{1}}+Y_{\eta_{R}}$ converges to $Y_{N_{1}}^{\infty}$ at $z_{2}=300$ in the case $M_{1}=1.5$ TeV. These examples show that $N_{1}$ in a wide mass range can explain the required relic abundance of DM in a consistent way with successful leptogenesis by tuning $h_{1}$ and $\lambda_{3,4}$ suitably. Couplings $\lambda_{3}$ and $\lambda_{4}$ play a crucial role especially when $M_{1}$ and $M_{\eta_{R}}$ take same order values since $Y_{N_{1}}^{\infty}~{}{{}^{>}_{\sim}}~{}Y_{\eta_{R}}^{\infty}$ should be satisfied in that case. We find that the lightest right-handed neutrino $N_{1}$ can be a good DM candidate consistently with leptogenesis and the model presents an interesting possibility such that three unsolved problems in the SM can be explained by low scale right-handed neutrinos simultaneously.
Finally, we should address phenomenological signatures of the model. Extra charged and neutral scalars $\eta^{\pm}$ and $\eta_{R,I}$ might be directly examined in the future collider experiments [31]. However, if $N_{1}$ has a very small $h_{1}$, it might be difficult to distinguish the model from the one in which DM is $\eta_{R}$ through collider signatures. The existence of right-handed neutrinos and the additional inert doublet scalar at TeV scales are expected to cause a new non-negligible contribution at one-loop level to LFV processes like $\mu\rightarrow e\gamma$ and also the electric dipole moment of electron (EDME). It is useful to examine whether their present experimental bounds are satisfied and also to give their prospect in the future experiments. If we use the neutrino Yukawa couplings fixed by eqs. (7) and (12), we can estimate them definitely. Branching ratio of the LFV process $\ell_{\alpha}\rightarrow\ell_{\beta}\gamma$ in the scotogenic model is expressed as [32]
$B(\ell_{\alpha}\rightarrow\ell_{\beta}\gamma)=\frac{3\alpha}{64\pi(G_{F}M_{\eta}^{2})^{2}}\left|\sum_{k=1}^{3}U_{\alpha k}^{\ast}U_{\beta k}h_{k}^{2}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\right|^{2},$ (28)
and the EDME is estimated as [14]
$d_{e}/e=\sum_{k=1}^{3}\frac{U_{ek}^{\ast}U_{ek}h_{k}^{2}m_{e}}{16\pi^{2}M_{\eta}^{2}}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\sin\gamma_{k},$ (29)
where $F_{2}(x)$ is defined as
$F_{2}(x)=\frac{1-6x^{2}+3x^{4}+2x^{6}-6x^{4}\ln x^{2}}{6(1-x^{2})^{4}}.$ (30)
Although these could be possible signatures of the model, they are respectively of $O(10^{-19})$ and $O(10^{-35})$ cm for the parameters given in eq. (14) which are used to explain both the DM relic abundance and the baryon number asymmetry in the present study. These values are much smaller than the present bounds [19], and they seem to be difficult to be reached in near future experiments. These results are not expected to be changed largely by varying the assumed parameters.
4 Summary
Both existences of neutrino mass and DM are unexplained problems in the SM. If they are solved in a common framework, it could be a promising candidate for new physics. The scotogenic model is such an example. The lightest right-handed neutrino is one of the DM candidates in this model. Although it could have a distinguishable character such that it cannot be found through direct detection experiments, it is known to be difficult to satisfy both constraints from the relic abundance and the LFV bounds simultaneously since they require incompatible magnitude for its Yukawa couplings. In order to escape this problem, additional interaction like $U(1)^{\prime}$ has been introduced to the right-handed neutrinos usually. Since it makes them interact with nuclei and then the direct detection experiments are relevant, the nature distinct from other ordinary WIMP candidates seems to be lost. In the present situation where the direct detection experiments exclude a lot of DM candidates, it may be an interesting possibility that DM is a right-handed neutrino which has no other interaction than the neutrino Yukawa couplings.
In this study, we propose a scenario in the original scotogenic model which allows the lightest right-handed neutrino can be DM without extra interaction. We show that the DM abundance and the baryon number asymmetry can be simultaneously explained if we assume certain mass degeneracy among new ingredients of the model. Although strict mass degeneracy is required, it is noticeable that three unsolved problems in the SM can be explained through the right-handed neutrinos simultaneously. An interesting point of this scenario is that the relic abundance of DM is determined not only by the interaction of DM itself but also by the interaction of its mother particle $\eta_{R}$. The model gives an example that the interaction searched by the direct detection experiments of DM is irrelevant to the interaction which determines the DM abundance.

### v2

@@
 A right-handed neutrino is a promising candidate for dark matter (DM) which has no interaction with nuclei.
-Since two right-handed neutrinos explain neutrino oscillation data and baryon number asymmetry through seesaw mechanism and leptogenesis based on the out-of-equilibrium decay of the lighter one, the lightest third right-handed neutrino could be free from them and stable to be DM.
+Since two right-handed neutrinos explain neutrino oscillation data and baryon number asymmetry through both the seesaw mechanism and leptogenesis based on the out-of-equilibrium decay of the lighter one, the lightest third right-handed neutrino could be free from them and stable to be DM.
 It suggests a possibility that the right-handed neutrinos could be common key particles for the explanation of three problems in the standard model (SM).
@@
 Neutrino oscillation experiments prove the existence of nonzero neutrino mass [1].
-Observations of the abundances of light elements and the cosmic microwave background suggest the existence of the baryon number asymmetry in the Universe [2, 3].
+Observations of both the abundances of light elements and the cosmic microwave background suggest the existence of baryon number asymmetry in the Universe [2, 3].
 However, both the neutrino mass and the baryon number asymmetry cannot be explained in the standard model (SM).
@@
 On the other hand, the SM has no candidate of dark matter (DM) whose existence has been proved through various astrophysical and cosmological observations.
-Direct detection experiments of it severely constrain its mass and scattering cross section with a nucleus, and they exclude a lot of DM candidates from a consistency with the required DM relic abundance [7].
+Its direct detection experiments severely constrain its mass and elastic scattering cross section with a nucleus.
+As a result, a lot of DM candidates are excluded from their consistency with the required DM relic abundance [7].
 It may suggest a possibility that DM has no interaction with nuclei.
 In this sense, the right-handed neutrino, which is massive and electrically neutral, could be a good candidate of DM since it has no interaction with nuclei.
+A possibility such that a right-handed neutrino is DM has been studied from a various view points.
+One may find recent works in [8, 9, 10, 11].
 If we combine this idea with the above solution for the neutrino mass and the baryon number asymmetry, the third right-handed neutrino seems to be a good DM candidate if its stability is guaranteed by some symmetry.
-This stability requirement may impose to introduce an additional doublet scalar $\eta$ since we have to guarantee the existence of their invariant Yukawa couplings between doublet leptons and right-handed neutrinos under the assumed symmetry.
+This requirement for the symmetry may impose to introduce an additional doublet scalar $\eta$ since we have to guarantee the existence of their invariant Yukawa couplings between doublet leptons and right-handed neutrinos under the assumed symmetry.
 As such an example, we have the scotogenic model [12], which can make the neutrino mass generation possible radiatively around TeV scales.
@@
 In previous studies aiming to solve these problems including the case where DM is the lightest neutral component of $\eta$, the model is extended by some way, for example, new interaction is introduced to the right-handed neutrinos [15, 16, 17, 18].
-In this paper, we propose a solution to these problems within the original scotogenic model without considering any extension of the model.aaaIt is discussed in [15] that $N_{1}$ can be DM consistently with the LFV constraint by considering coannihilation among the right-handed neutrinos without any extension.
+In this paper, we propose a solution to these problems within the original scotogenic model without considering any extension of the model.aaaIt is discussed that $N_{1}$ can be DM consistently with the LFV constraint by considering coannihilation among the right-handed neutrinos without any extension [19].
 The remaining parts are organized as follows.
-In section 2, we review the model and discuss a neutrino mass matrix derived in it to fix flavor mixing and mass eigenvalues.
-After that, we study a consistent solution with both the relic DM abundance and the constraint from the lepton flavor violating processes.
+In section 2, after we review the model briefly, we discuss a neutrino mass matrix derived from it in detail, and flavor mixing and mass eigenvalues are fixed.
+Using them, we study a solution consistent with both the relic DM abundance and the constraint from the lepton flavor violating processes.
 In section 3, we discuss the consistency of the scenario with successful leptogenesis.
@@
 for the vacuum stability.
-If $\eta$ is assumed to have no vacuum expectation value (VEV), neutrino mass is not produced at tree level and the $Z_{2}$ symmetry is kept as an exact asymmetry of the model.
+If $\eta$ is assumed to have no vacuum expectation value (VEV), neutrino mass is not produced at tree level and the $Z_{2}$ symmetry is kept as an exact symmetry of the model.
 After the electroweak symmetry breaking, mass of the components of $\eta$ is fixed through the Higgs VEV as
 $M_{\eta^{\pm}}^{2}=m_{\eta}^{2}+\lambda_{3}\langle\phi\rangle^{2},\quad M_{\eta_{R}}^{2}=m_{\eta}^{2}+(\lambda_{3}+\lambda_{4}+\lambda_{5})\langle\phi\rangle^{2},\quad M_{\eta_{I}}^{2}=m_{\eta}^{2}+(\lambda_{3}+\lambda_{4}-\lambda_{5})\langle\phi\rangle^{2},$ (3)
-where $\eta^{\pm}$ stands for charged components, and $\eta_{R}$ and $\eta_{I}$ are real and imaginary components, respectively.
+where $\eta^{\pm}$ stands for charged components.
+$\eta_{R}$ and $\eta_{I}$ are real and imaginary part of the neutral component, respectively.
 In this setting, a lightest right-handed neutrino $N_{1}$ or $\eta_{R}$ can be stable due to the $Z_{2}$ symmetry.
@@
 2.1 Neutrino mass and flavor mixing
-Although neutrino mass is forbidden at tree level in Lagrangian (1), small neutrino mass is induced through a one-loop diagram caused by the Higgs VEV as
+Although neutrino mass is forbidden at tree level in Lagrangian (1), small neutrino mass is induced through a one-loop diagram caused by the Higgs VEV $\langle\phi\rangle$ as
 $({\cal M}_{\nu})_{\alpha\beta}=\sum_{k}h_{\alpha k}h_{\beta k}\Lambda_{k}e^{i\gamma_{k}}\equiv\sum_{k}\mu_{\alpha\beta}^{k}\Lambda_{k}e^{i\gamma_{k}},$
@@
 Its diagonalization matrix can be described by using the Yukawa couplings $h_{\alpha k}$ explicitly.
-When a condition $\sum_{\alpha}(h_{\alpha j}h_{\alpha k})\propto\delta_{jk}$ is satisfied, which corresponds to the condition that matrices $\mu^{j}$ and $\mu^{k}$ are commutable [11], it is diagonalized by a matrix $U$ as $U^{T}{\cal M}_{\nu}U={\rm diag}(m_{1},m_{2},m_{3})$.
+When a condition $\sum_{\alpha}(h_{\alpha j}h_{\alpha k})\propto\delta_{jk}$ is satisfied, which corresponds to the condition that matrices $\mu^{j}$ and $\mu^{k}$ in eq.
+(5) are commutable [15], ${\cal M}_{\nu}$ is diagonalized by a matrix $U$ as $U^{T}{\cal M}_{\nu}U={\rm diag}(m_{1},m_{2},m_{3})$.
 This $U$ can be expressed as
 $U=(U_{1},U_{2},U_{3}),\quad U_{k}^{T}\equiv(U_{ek},U_{\mu k},U_{\tau k}),\quad U_{\alpha k}\equiv\frac{h_{\alpha k}}{h_{k}},$ (6)
-where $h_{k}$ is a real constant. The absolute value of mass eigenvalues $m_{k}$ is fixed as
+where $h_{k}$ is a real constant. An absolute value of the mass eigenvalue $m_{k}$ is fixed as
 $m_{k}=\left|\sum_{\alpha}h_{\alpha k}^{2}\right|\Lambda_{k},=p_{k}h_{k}^{2}\Lambda_{k},$ (7)
 where $p_{k}$ is defined as $p_{k}=\left|\sum_{\alpha}U_{\alpha k}^{2}\right|$.
-Here, it should be noted that $\sum_{\alpha}U_{\alpha k}^{2}=1$ is not satisfied if Yukawa couplings $h_{\alpha k}$ are complex.
-By using this $U$, the PMNS matrix [17] can be written as $V_{\rm PMNS}=U_{e}^{\dagger}UP$ , where $U_{e}$ is a matrix which is used to diagonalize the charged lepton mass matrix and $P$ is a Majorana phase matrix ${\rm diag}(1,e^{i(\gamma_{2}-\gamma_{1})},e^{i(\gamma_{3}-\gamma_{1})})$.
+Here, it should be noted that $\sum_{\alpha}U_{\alpha k}^{2}=1$ is not satisfied if the Yukawa coupling $h_{\alpha k}$ is complex.
+By using this $U$, the PMNS matrix [21] can be written as $V_{\rm PMNS}=U_{e}^{\dagger}UP$ , where $U_{e}$ is a matrix which is used to diagonalize a charged lepton mass matrix and $P$ is a matrix of Majorana phases ${\rm diag}(1,e^{i(\gamma_{2}-\gamma_{1})},e^{i(\gamma_{3}-\gamma_{1})})$.
 Although $V_{\rm PMNS}$ is unitary in a case where $h_{\alpha k}$ is real, it is not unitary in a case where $h_{\alpha k}$ is complex since $U$ is not a unitary matrix.
 Since $U$ is the matrix which diagonalizes a symmetric matrix, it may be parametrized as
-$U_{\nu}=\left(\begin{array}[]{ccc}c_{12}c_{13}&s_{12}c_{13}&s_{13}\\ -s_{12}c_{23}-c_{12}s_{13}s_{23}&c_{12}c_{23}-s_{12}s_{13}s_{23}&c_{13}s_{23}\\ s_{12}s_{23}-c_{12}s_{13}c_{23}&-c_{12}s_{23}-s_{12}s_{13}c_{23}&c_{13}c_{23}\\ \end{array}\right),$ (8)
+$U=\left(\begin{array}[]{ccc}c_{12}c_{13}&s_{12}c_{13}&s_{13}\\ -s_{12}c_{23}-c_{12}s_{13}s_{23}&c_{12}c_{23}-s_{12}s_{13}s_{23}&c_{13}s_{23}\\ s_{12}s_{23}-c_{12}s_{13}c_{23}&-c_{12}s_{23}-s_{12}s_{13}c_{23}&c_{13}c_{23}\\ \end{array}\right),$ (8)
 where $c_{ij}=\cos\theta_{ij}$ and $s_{ij}=\sin\theta_{ij}e^{i\alpha_{ij}}$.
-We can represent the neutrino Yukawa couplings $h_{\alpha k}$ and mass eigenvalues $m_{k}$ by using this $U$ as eq.
+We can represent the neutrino Yukawa coupling $h_{\alpha k}$ and the mass eigenvalue $m_{k}$ by using this $U$ as eq.
 (6) and eq.
 (7), respectively.
-If $h_{\alpha 1}$ has the same structure as $h_{\alpha 2}$ for example, the mass matrix can be diagonalized by $U$ given in eq.
-(8) but we have a zero mass eigenvalue $m_{1}=0$.
-On the $m_{2}$ formula in eq.
+If $h_{\alpha 1}$ has the same structure as $h_{\alpha 2}$ for example, the mass matrix ${\cal M}_{\nu}$ can be diagonalized by $U$ given in eq.
+(8) but it has a zero mass eigenvalue $m_{1}=0$.
+On the formula of $m_{2}$ in eq.
 (7), $p_{2}h_{2}^{2}\Lambda_{2}$ is replaced by $p_{1}h_{1}^{2}\Lambda_{1}+p_{2}h_{2}^{2}\Lambda_{2}$.
@@
 $h_{ei}=h_{\mu i}=-h_{\tau i}=\frac{1}{\sqrt{3}}h_{i}~~(i=1,2),\quad h_{e3}=0,~h_{\mu 3}=h_{\tau 3}=\frac{h_{3}}{\sqrt{2}}.$ (9)
-The mixing matrix $U$ can be taken as the tribimaximal mixing matrix [18] by fixing the mixing angles as
+In that case, $U$ is found to be the tribimaximal mixing matrix [22] and the mixing angles are fixed as
 $s_{12}=\frac{1}{\sqrt{3}},\quad s_{23}=\frac{1}{\sqrt{2}},\quad s_{13}=0,\quad\alpha_{12}=\alpha_{23}=0,$ (10)
@@
 $m_{1}=0,\quad m_{2}=h_{1}^{2}\Lambda_{1}+h_{2}^{2}\Lambda_{2},\quad m_{3}=h_{3}^{2}\Lambda_{3}.$ (11)
-Since the fault of the tribimaximal mixing matrix as the PMNS matrix could be overcome by considering a nontrivial flavor mixing matrix $U_{e}$ in the charged lepton sector [14], it can become a phenomenologically interesting possibility.
+The fault of the tribimaximal mixing matrix as the PMNS matrix could be overcome by considering a nontrivial flavor mixing matrix $U_{e}$ in the charged lepton sector [18].
+Since $U_{e}$ is expected to be almost diagonal, it can become a phenomenologically viable possibility.
+Here, we remark on an important consequence caused by the feature of this mass matrix addressed above.
+If we consider leptogenesis based on the out-of-equilibrium decay of the right-handed neutrino $N_{k}$, $CP$ asymmetry in this process is known to be caused by the interference between tree and one-loop diagrams and it should be nonzero at least to make leptogenesis work.
+On this point, the model predicts an interesting result.
+Since this $CP$ asymmetry is proportional to ${\rm Im}(\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j})^{2}$ for $k\not=j$ as shown in eqs.
+(21) and (25) later, this condition cannot be satisfied because of $\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j}\propto\delta_{kj}$ for the real neutrino Yukawa coupling $h_{\alpha k}$.
+This is the case even if Majorana phases in the right-handed neutrino mass are taken into account.
+Thus, the Yukawa coupling $h_{\alpha k}$ should be complex, and the PMNS matrix cannot be unitary in that case.
+However, we should note that there is an exceptional case where the $CP$ asymmetry can take a nonzero value even if the Yukawa coupling $h_{\alpha k}$ is real.
+The existence of a zero mass eigenvalue presents such a case as found in an example given in eq.
+(9) when Majorana phases exist in the right-handed neutrino mass [18].
+The PMNS matrix becomes unitary in this case.
 Neutrino oscillation data fix squared mass differences $\Delta m_{32}^{2}\equiv m_{3}^{2}-m_{2}^{2}$ and $\Delta m_{21}^{2}\equiv m_{2}^{2}-m_{1}^{2}$.
@@
 $m_{2}=\sqrt{\Delta m_{21}^{2}},\qquad m_{3}=\sqrt{\Delta m_{32}^{2}+\Delta m_{21}^{2}}.$ (12)
-If we use the best fit values from neutrino oscillation data given in [19] for $\sin\theta_{ij}$ in the matrix $U$, we can calculate two effective masses relevant to two observables, that is, the effective masses for the neutrinoless double $\beta$ decay and the $\beta$ decay.
+If we use the best fit values from neutrino oscillation data given in [23] for $\Delta m_{ij}^{2}$ and $\sin\theta_{ij}$ in the matrix $U$, we can calculate two effective masses relevant to two observables, that is, the effective masses for the neutrinoless double $\beta$ decay and the $\beta$ decay.
 They are predicted to be
-$m_{\beta\beta}\simeq\left|\sum_{j=1}^{3}U_{ej}^{2}m_{j}\right|=0.0037~{}{\rm eV},\quad m_{\beta}\simeq\sqrt{\sum_{j=1}^{3}|U_{ej}|^{2}m_{j}^{2}}=0.0088~{}{\rm eV}.$ (13)
-These results are much smaller than their present bounds [20, 21].
+$m_{\beta\beta}\simeq\left|\sum_{j=1}^{3}U_{ej}^{2}m_{j}\right|=0.0037~{\rm eV},\quad m_{\beta}\simeq\sqrt{\sum_{j=1}^{3}|U_{ej}|^{2}m_{j}^{2}}=0.0088~{\rm eV}$ (13)
+in normal ordering case. These results are much smaller than their present bounds [24, 25].
 […]
-$M_{1}=1.5~{}{\rm TeV},\quad M_{\eta_{R}}\simeq M_{2}\simeq M_{3}\simeq 3~{}{\rm TeV},\quad\lambda_{5}=-10^{-3},$
-$h_{1}=6\times 10^{-9},\quad h_{2}=1.11\times 10^{-3},\quad h_{3}=4.15\times 10^{-3},$ (14)
-where we assume $\alpha_{12}=\alpha_{13}=0,~{}\alpha_{23}=\frac{\pi}{2}$ and use the best fit values from neutrino oscillation data for $\sin\theta_{ij}$.
-A smaller $|\lambda_{5}|$ makes $h_{k}$ larger for fixed values of $M_{k}$.
-Since DM is supposed to be $N_{1}$ here, there is no constraint on a lower bound of the value of $|\lambda_{5}|$ differently from the case where $\eta_{R}$ is DM [10].
-We find that the present bounds of LFV like $\mu\rightarrow e\gamma$ [19] can be satisfied for these parameters.
-We come back to this point later.
-In the DM study, it may be useful to note that the above values of $h_{2,3}$ are not affected even if $h_{1}$ takes the same order value as $h_{2,3}$ as long as $M_{1}$ is much smaller than $M_{\eta_{R}}$.
-It is caused by the fact that $\Lambda_{1}$ has a suppression of $O(M_{1}/M_{\eta_{R}})$ compared with $\Lambda_{2,3}$.
-It allows $N_{1}$ with keV scale mass to have an acceptable relic abundance as DM for the same $h_{2,3}$ given in eq.
-(14).
-Finally, we note an important consequence caused by the feature of this mass matrix addressed above.
-If we consider leptogenesis based on the out-of-equilibrium decay of the right-handed neutrino $N_{k}$, the model predicts an interesting result.
-$CP$ asymmetry in this process is known to be caused by the interference between tree and one-loop diagrams, and it should be nonzero at least to make leptogenesis work.
-Since it is proportional to ${\rm Im}(\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j})^{2}$ for $k\not=j$ as shown in eqs.
-(21) and (25) later, this condition cannot be satisfied because of $\sum_{\alpha}h_{\alpha k}^{\ast}h_{\alpha j}\propto\delta_{kj}$ for real neutrino Yukawa couplings $h_{\alpha k}$ even if Majorana phases in the right-handed neutrino mass are taken into account.
-Thus, Yukawa couplings $h_{\alpha k}$ should be complex, and the PMNS matrix cannot be unitary in that case.
-However, we should note that there is an exceptional case where the $CP$ asymmetry can take a nonzero value even if Yukawa couplings $h_{\alpha k}$ are real.
-The existence of a zero mass eigenvalue presents such a case as found in an example given in eq.
-(9) when Majorana phases exist in the right-handed neutrino mass [14].
-The PMNS matrix becomes unitary in this case.
+$M_{\eta_{R}}\simeq M_{2}\simeq M_{3}\simeq 3~{\rm TeV},\quad\lambda_{5}=-10^{-3},$
+$h_{2}=1.11\times 10^{-3},\quad h_{3}=4.15\times 10^{-3},$ (14)
+where we assume $\alpha_{12}=\alpha_{13}=0,~\alpha_{23}=\frac{\pi}{2}$ and use for $\sin\theta_{ij}$ the best fit values from neutrino oscillation data.
+$M_{1}$ and $h_{1}$ can be fixed freely as long as $h_{1}$ is small enough.
+We note that the present bounds of LFV like $\mu\rightarrow e\gamma$ [23] can be satisfied for these parameters.
+We come back to this point in the last part of the paper again.
 2.2 Relic abundance of right-handed neutrino DM
@@
 However, its abundance is difficult to reach the required value in the freezeout scenario as mentioned in the introduction.
-It can reach the equilibrium at the temperature higher than its mass through both the $\eta$ decay and the scatterings caused by the neutrino Yukawa couplings $h_{\alpha 1}$ if they have sufficient magnitude.
-In that case, $h_{\alpha 1}$ has to take a large value of $O(1)$ to reduce its abundance sufficiently through the annihilation.
+In fact, it can be produced in the thermal bath at the temperature higher than its mass through both the $\eta$ decay and the scatterings ($\eta^{\dagger}\eta\rightarrow N_{1}N_{1}^{c}$ and $\ell_{\alpha}^{\dagger}\ell_{\alpha}\rightarrow N_{1}N_{1}^{c}$) caused by the neutrino Yukawa coupling $h_{\alpha 1}$ if it is large enough.
+In that case, $h_{\alpha 1}$ has to take a value of $O(1)$ to reduce its abundance sufficiently through the annihilation.
 Unfortunately, such a large Yukawa coupling causes LFV processes like $\mu\rightarrow e\gamma$ at an unacceptable level through the one-loop diagram [13].
@@
 (14), its thermal abundance is originally difficult to reach a value larger than the required one so that the freezeout scenario cannot be applied to the model.
-However, even in that case, we can consider an alternative possibility such that it is realized through the decay of $\eta_{R}$ following the freezein scenario [22].
-$N_{1}$ is produced through the $\eta_{R}$ decay caused by the neutrino Yukawa coupling $h_{1}$.
-Since $h_{1}$ is supposed to be sufficiently small here, the $N_{1}$ generation through the $\eta_{R}$ decay is expected to be largely suppressed.
-On the other hand, the next lightest DM candidate $\eta_{R}$ is in the thermal equilibrium through the SM interactions.
-Decrease of the number density of $\eta_{R}$ is controlled by both the coannihilation [23] to the SM contents and the decay to $N_{1}$.
-The $N_{1}$ relic abundance is determined by both the $N_{1}$ freezein and the $\eta_{R}$ freezeout depending on their relative effectiveness.
+However, even in that case, we can consider the freezein scenario [26] as an alternative possibilitydddA right-handed neutrino DM through the freezein mechanism has been studied in the different model [8] and in the original scotogenic model [10, 11].
+in which $N_{1}$ is produced through the $\eta$ decay and the scatterings caused by the neutrino Yukawa coupling $h_{1}$.
+Since $h_{1}$ is supposed to be sufficiently small here, the $N_{1}$ generation through these processes is expected to be largely suppressed.
+On the other hand, the next lightest DM candidate $\eta_{R}$ is in the thermal equilibrium and decrease of its number density is controlled by its coannihilation [27] caused by the SM interactions.
+After the freezeout of this coannihilation, the relic $\eta_{R}$ finally decays to $N_{1}$.
+Thus, we have to take into account that there is a case where the contribution from such $N_{1}$ plays a crucial role in the estimation of the $N_{1}$ relic abundance.
 We express the number density of $\eta_{R}$ and $N_{1}$ in the comoving volume by $Y_{i}=\frac{n_{i}}{s}$, where $n_{i}$ is the number density of particle species $i$ and $s$ is the entropy density.
@@
 where $H(M_{2})$ represents the Hubble parameter at $T=M_{2}$.
-Reaction densities $\gamma^{D}(a)$ and $\gamma^{S}(ab)$ are respectively related to thermally averaged decay width $\langle\Gamma_{a}\rangle$ for $a\rightarrow ij$ and cross section $\langle\sigma_{ab}|v|\rangle$ for scattering $ab\rightarrow ij$ as[25]
+Reaction densities $\gamma^{D}(a)$ and $\gamma^{S}(ab)$ are respectively related to thermally averaged decay width $\langle\Gamma_{a}\rangle$ for $a\rightarrow ij$ and thermally averaged cross section $\langle\sigma_{ab}|v|\rangle$ for the scattering $ab\rightarrow ij$ as[29]
 $\gamma^{D}(a)=\langle\Gamma_{a}\rangle n_{a}^{\rm eq},\qquad\gamma^{S}(ab)=\langle\sigma_{ab}|v|\rangle n_{a}^{\rm eq}n_{b}^{\rm eq},$ (17)
@@
 The present DM abundance can be expressed as $\Omega_{DM}=M_{DM}Y_{DM}^{\infty}s_{0}/\rho_{0}$, where $\rho_{0}$ and $s_{0}$ are the energy density and the entropy density in the present Universe.
-They are given as $\rho_{0}=3M_{\rm pl}^{2}H_{0}^{2}$ and $s_{0}=2.27\times 10^{-38}$ GeV3, and then it determines the required $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ through $\Omega_{N_{1}}h^{2}=0.12$ as $Y_{N_{1}}^{\infty}=4.2\times 10^{-10}(M_{1}/{\rm GeV})^{-1}$.
-Here, we roughly estimate both the freezeout temperature $T_{\eta}$ of the coannihilation of $\eta_{R}$ and a value of $h_{1}$ for the $\eta_{R}$ decay to reach the equilibrium at $T_{\eta}$.
-The freezeout temperature $T_{\eta}(\equiv M_{\eta_{R}}/z_{\eta})$ can be estimated through a condition $2n_{\eta_{R}}^{\rm eq}(T_{\eta})\langle\sigma_{\eta\eta}|v|\rangle=H(T_{\eta})z_{\eta}$ by assuming instantaneous freezeout as found from eq.
-(15).
-Using the Hubble parameter $H(T)$ given in eq.
-(19), $z_{\eta}$ can be determined by solving this condition as [26]
-$z_{\eta}=\ln\frac{0.384M_{\rm pl}\langle\sigma_{\eta\eta}|v|\rangle_{\eta}M_{\eta_{R}}}{(g_{\ast}z_{\eta})^{1/2}}.$ (18)
-The $h_{1}$ for the $\eta_{R}$ decay to reach the equilibrium at $T_{\eta}$ can be estimated from the condition $\Gamma_{\eta_{R}}=H(T_{\eta})$.
+Since they are given as $\rho_{0}=3M_{\rm pl}^{2}H_{0}^{2}$ and $s_{0}=2.27\times 10^{-38}$ GeV3, the required $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is found to be expressed as $Y_{N_{1}}^{\infty}=4.2\times 10^{-10}(M_{1}/{\rm GeV})^{-1}$ from $\Omega_{N_{1}}h^{2}=0.12$.
+Here, we roughly estimate both the freezeout temperature $T_{\eta}$ of the coannihilation of $\eta_{R}$ and a value of $h_{1}$ which makes the $\eta_{R}$ decay reach the equilibrium at $T_{\eta}$.
+Such a value of $h_{1}$ can be estimated from the condition $\Gamma_{\eta_{R}}=H(T_{\eta})$.
 The $\eta_{R}$ decay width $\Gamma_{\eta_{R}}$ and the Hubble parameter $H(T)$ are given as
@@
 $h_{1}c_{p}=3\times 10^{-8}\left(\frac{M_{\eta_{R}}}{3~{\rm TeV}}\right)^{1/2},$ (19)
-where $c_{p}$ is a constant determined by the phase space volume.cccWhen the mass of $\eta_{R}$ and $N_{1}$ take closely degenerate values, $c_{p}$ can be represented as $c_{p}=2\delta_{1}$, where $\delta_{1}$ is defined by $M_{1}=(1-\delta_{1})M_{\eta_{R}}$ with $\delta_{1}\ll 1$.
+where $c_{p}$ is a constant determined by the phase space volume.eeeWhen the masses of $\eta_{R}$ and $N_{1}$ take closely degenerate values, $c_{p}$ can be represented as $c_{p}=2\delta_{1}$ where $\delta_{1}$ is defined by $M_{1}=(1-\delta_{1})M_{\eta_{R}}$ with $\delta_{1}\ll 1$.
+The freezeout temperature $T_{\eta}(\equiv M_{\eta_{R}}/z_{\eta})$ can be estimated through a condition $2n_{\eta_{R}}^{\rm eq}(T_{\eta})\langle\sigma_{\eta\eta}|v|\rangle=H(T_{\eta})z_{\eta}$ as found from eq.
+(15) by assuming instantaneous freezeout.
+Using the Hubble parameter $H(T)$ given in eq.
+(18), $z_{\eta}$ can be determined by solving this condition as [30]
+$z_{\eta}=\ln\frac{0.384M_{\rm pl}\langle\sigma_{\eta\eta}|v|\rangle M_{\eta_{R}}}{(g_{\ast}z_{\eta})^{1/2}}.$ (20)
 If $h_{1}$ takes a larger value than the one given in eq.
-(20), the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be approximated well as $Y_{N_{1}}^{\infty}=Y_{\eta_{R}}^{\rm eq}(T_{c})$ through the freezein caused by the $\eta_{R}$ decay, where $T_{c}$ is defined by $Y_{N_{1}}(T_{c})=Y_{\eta_{R}}^{\rm eq}(T_{c})$ since the $\eta_{R}$ decay at $T<T_{c}$ does not affect $Y_{N_{1}}^{\infty}$ substantially.
-$Y_{N_{1}}^{\infty}$ is determined by the coupling $h_{1}$ for a fixed $M_{1}$ and the freezeout of the $\eta_{R}$ (co)annihilation does not affect the result.
+(19), the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be approximated well as $Y_{N_{1}}^{\infty}=Y_{\eta_{R}}^{\rm eq}(T_{c})$ through the freezein caused by the $\eta_{R}$ decay.
+$T_{c}$ is defined by $Y_{N_{1}}(T_{c})=Y_{\eta_{R}}^{\rm eq}(T_{c})$ since the $\eta_{R}$ decay at $T<T_{c}$ does not affect $Y_{N_{1}}^{\infty}$ substantially.
+$Y_{N_{1}}^{\infty}$ is determined only by the coupling $h_{1}$ for a fixed $M_{1}$ and the freezeout of the $\eta_{R}$ coannihilation does not affect this result.
 However, this situation could change when $h_{1}$ takes a smaller value than the one given in eq.
 (19) and $T_{c}<T_{\eta}$ is satisfied.
-In this case, the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be affected by the equilibrium number density of $\eta_{R}$ at $T_{\eta}$.
-$T_{\eta}$ has been known to be determined through eq.
-(18) by the couplings $\lambda_{3}$ and $\lambda_{4}$ through the study of the inert doublet DM model [24].
-If we use this knowledge, the required $N_{1}$ relic abundance could be predicted through a relation $Y_{N_{1}}^{\infty}\simeq Y_{\eta_{R}}^{\rm eq}(T_{\eta})$.
+In this case, the $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$ is expected to be affected by the equilibrium number density of $\eta_{R}$ at $T_{\eta}$ since $\eta_{R}$ decays to $N_{1}$ finally.
+$T_{\eta}$ is determined through eq.
+(20) by $\langle\sigma_{\eta\eta}|v|\rangle$ which depends on the couplings $\lambda_{3}$ and $\lambda_{4}$.
+In the study of the inert doublet DM model [28], these couplings are known to play a crucial role there.
+If we take account of it, the $N_{1}$ relic abundance could be predicted through a relation $Y_{N_{1}}^{\infty}\simeq Y_{\eta_{R}}^{\rm eq}(T_{\eta})$.
 It also gives constraints on the couplings $\lambda_{3}$ and $\lambda_{4}$.
 Fig.
-1   Left panel: Evolution of the number density of $N_{1}$ and $\eta_{R}$ in the comoving volume for three values of $M_{1}$ are plotted with red lines, 1.5 TeV (solid), 150 GeV (dashed) and 1.5 GeV (dotted).
+1   Left panel: Evolution of the number density of $N_{1}$ and $\eta_{R}$ in the comoving volume.
+They are plotted with red lines for three values of $M_{1}$, 1.5 TeV (solid), 150 GeV (dashed) and 1.5 GeV (dotted).
 $h_{1}$ is fixed at $7.8\times 10^{-9}$, $4.5\times 10^{-8}$ and $7.0\times 10^{-6}$ for each $M_{1}$, respectively.
 $M_{\eta_{R}}=3$ TeV and $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$ are assumed.
-Horizontal black lines represent $Y_{N_{1}}$ values required for each value of $M_{1}$ to realize the present DM relic density.
-While an $Y_{\eta_{R}}$ line for $M_{1}=1.5$ TeV deviates from $Y_{\eta_{R}}^{\rm eq}$, $Y_{\eta_{R}}$ in other two cases coincides with $Y_{\eta_{R}}^{\rm eq}$ since the $\eta_{R}$ decay is already in the equilibrium at a temperature higher than $T_{c}$.
+Horizontal black lines represent $Y_{N_{1}}^{\infty}$ values required to realize the present DM relic density for each value of $M_{1}$.
+While $Y_{\eta_{R}}$ for $M_{1}=1.5$ TeV deviates from $Y_{\eta_{R}}^{\rm eq}$, $Y_{\eta_{R}}$ in other two cases coincides with $Y_{\eta_{R}}^{\rm eq}$ since the $\eta_{R}$ decay is already in the equilibrium at a temperature higher than $T_{\eta}$.
 Right panel: Contours of $Y_{\eta_{R}}(T_{\eta})$ for $M_{\eta_{R}}=3$ TeV in the $(\lambda_{3},\lambda_{4})$ plane.
@@
 (14), we plot the solutions in the left panel for the cases $(M_{1},h_{1})=(1.5~{\rm TeV},7.8\times 10^{-9})$, $(150~{\rm GeV},4.5\times 10^{-8})$, and $(1.5~{\rm GeV},7.0\times 10^{-6})$ where $h_{1}$ is tuned to give the required DM abundance.
-In this panel, $Y_{N_{1}}$ required to explain the DM abundance is shown by horizontal black lines for each case.
-In the latter two cases, the required $Y_{N_{1}}^{\infty}$ is realized through the pure freezein due to the $\eta_{R}$ decay.
-$Y_{\eta_{R}}$ does not deviate from the equilibrium one since its decay to $N_{1}$ is already in the equilibrium at $T>T_{\eta}$ and then $Y_{\eta_{R}}=Y_{\eta_{R}}^{\rm eq}$ is kept.
+In this panel, $Y_{N_{1}}^{\infty}$ required to explain the DM abundance is shown by horizontal black lines for each case.
+In the latter two cases, $T_{c}>T_{\eta}$ is satisfied and the required $Y_{N_{1}}^{\infty}$ is found to be realized by the pure freezein due to the $\eta_{R}$ decay in the thermal equilibrium through a relation $Y_{N_{1}}^{\infty}=Y_{N_{1}}(T_{c})=Y_{\eta_{R}}^{\rm eq}(T_{c})$.
+$Y_{\eta_{R}}$ does not deviate from the equilibrium one there.
 The freezeout of the $\eta_{R}$ coannihilation in a case of $M_{\eta_{R}}=3$ TeV is shown in the right panel, in which contours for $Y_{\eta_{R}}^{\rm eq}(T_{\eta})=10^{-10.5},10^{-11},10^{-11.5},10^{-11.75}$, and $10^{-12}$ are plotted in the $(\lambda_{3},\lambda_{4})$ plane.
-Since $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$ is assumed in the left panel, we find $Y_{\eta_{R}}^{\rm eq}(T_{\eta})=10^{-11.75}$ is relevant.
-It is smaller than the required $Y_{N_{1}}^{\infty}$ in these cases, and then the freezeout of the $\eta_{R}$ coannihilation is irrelevant to the estimation of $Y_{N_{1}}^{\infty}$.
-As an extreme case of these examples, we may consider $N_{1}$ with keV scale mass which can be consistent with the neutrino oscillation data as discussed in the previous part.
+Since $(\lambda_{3},\lambda_{4})=(1.8,-1.0)$ is assumed in the left panel, $Y_{\eta_{R}}^{\rm eq}(T_{\eta})$ for it is smaller than the required $Y_{N_{1}}^{\infty}$ from the right panel.
+The freezeout of the $\eta_{R}$ coannihilation is irrelevant to the estimation of $Y_{N_{1}}^{\infty}$.
+As an extreme case for this kind of examples, we may consider $N_{1}$ with keV scale mass.
 We can confirm that the required relic DM abundance is obtained by tuning $h_{1}$ suitably, for example, $h_{1}=3.7\times 10^{-3}$ for $M_{1}=5$ keV.
-On the other hand, the freezeout of the $\eta_{R}$ coannihilation is relevant to the estimation of $Y_{N_{1}}$ in the case of $M_{1}=1.5$ TeV.
-Although $Y_{N_{1}}(T_{c})$ is smaller than the required value, the late time decay of the frozen $\eta_{R}$ contributes to $Y_{N_{1}}$ and make it the required value $Y_{N_{1}}^{\infty}$ as found from the figure.
+We note that the values of $h_{2,3}$ in eq.
+(14) are not affected even if $h_{1}$ takes the same order value as $h_{2,3}$ as long as $M_{1}$ is much smaller than $M_{\eta_{R}}$ as the present case.
+It is caused by the fact that $\Lambda_{1}$ has a strong suppression factor of $O(M_{1}/M_{\eta_{R}})$ compared with $\Lambda_{2,3}$.
+Thus, $N_{1}$ with keV scale mass is allowed to have an acceptable relic abundance as DM for $h_{2,3}$ given in eq.
+(14) without contradicting with the neutrino oscillation data.
+In the case of $M_{1}=1.5$ TeV, the freezeout of the $\eta_{R}$ coannihilation is relevant to the estimation of $Y_{N_{1}}$.
+Although $Y_{N_{1}}(T_{c})$ is smaller than the required value, the late time decay of the frozen $\eta_{R}$ contributes to $Y_{N_{1}}$ and makes it the required value $Y_{N_{1}}^{\infty}$ as found from the figure.
 Since the stability of the potential $V$ in eq.
 (1) requires $\lambda_{3,4}$ should satisfy the condition (2), an allowed region of $(\lambda_{3},\lambda_{4})$ is a part over the dashed line in the right panel.
-Although $Y_{N_{1}}$ is determined by the values of $\lambda_{3}$ and $\lambda_{4}$ in the $M_{1}=1.5$ TeV case, we note that there is no constraint on them from direct detection experiments since $N_{1}$ is DM which cannot interact with nuclei.
+If $\lambda_{3}$ and $\lambda_{4}$ are fixed to suitable values in this region, the required $Y_{N_{1}}^{\infty}$ can be obtained.
+Here, we should note that there is no constraint on the values of $\lambda_{3}$ and $\lambda_{4}$ from direct detection experiments since $N_{1}$ cannot interact with nuclei.
 It is a crucial difference from the case where $\eta_{R}$ is DM.
 The coupling constants $\lambda_{3}$ and $\lambda_{4}$ could be also constrained through the correction to the Higgs mass caused by them.
-This constraint is reported to be satisfied for $M_{\eta_{R}}\leq 4.4$ TeV in [27].
-In the above study, we do not take into account the effect caused by the simultaneous explanation of the baryon number asymmetry in the Universe through leptogenesis based on the $N_{2}$ decay.
-Since it could produce $\eta_{R}$ at late time, the $N_{1}$ abundance could be affected by it if the $N_{2}$ decay occurs near the freezeout of $\eta_{R}$.
+This constraint is reported in [31] to be satisfied for $M_{\eta_{R}}\leq 4.4$ TeV.
+In the above study, we do not take into account the effect caused by the simultaneous explanation of the baryon number asymmetry in the Universe through leptogenesis which is based on the $N_{2}$ decay.
+Since it could produce $\eta_{R}$ at late time, the $N_{1}$ abundance could be affected by it if the $N_{2}$ decay occurs near the freezeout of the coannihilation of $\eta_{R}$.
 We discuss this point in the next part.
@@
 In the present model, the baryon number asymmetry is considered to be produced through leptogenesis caused by the out-of-equilibrium decay of $N_{2}$.
-This requires that $N_{2}$ should be in the thermal bath and its decay to $\eta_{R}$ occurs in a $CP$ asymmetric way at a temperature higher than 100 GeV where sphaleron interaction is in the equilibrium.
+This requires that $N_{2}$ should be in the thermal bath and its decay to $\eta_{R}$ occurs in a $CP$ asymmetric way at a higher temperature than 100 GeV where sphaleron interaction is in the equilibrium.
 Values of $h_{2,3}$ given in eq.
-(14) constrained by the neutrino oscillation data are large enough to produce $N_{2,3}$ in the thermal bath as relativistic particles through the inverse decay and scatterings as long as reheating temperature is larger than $M_{3}$.
-$CP$ asymmetry in the decay $N_{2}\rightarrow\ell_{\alpha}\eta$ can be caused through an interference between a tree diagram and a one-loop diagram mediated by $N_{1}$ and $N_{3}$.
-The $CP$ asymmetry caused by one loop vertex diagrams is estimated through the formula
+(14) are large enough to produce $N_{2,3}$ in the thermal bath as relativistic particles through the inverse decay and scatterings caused by them as long as reheating temperature is larger than $M_{3}$.
+The $CP$ asymmetry in the decay $N_{2}\rightarrow\ell_{\alpha}\eta^{\dagger}$ can be caused through an interference between a tree diagram and a one-loop diagram mediated by $N_{1}$ and $N_{3}$.
+This $CP$ asymmetry is estimated through the formula
 $\varepsilon=\frac{1}{8\pi}\sum_{k=1,3}\frac{{\rm Im}\left[\left(\sum_{\alpha}h_{\alpha 2}h^{\ast}_{\alpha k}\right)^{2}\right]}{\sum_{\alpha}h_{\alpha 2}h_{\alpha 2}^{\ast}}F\left(\frac{M_{k}^{2}}{M_{2}^{2}},\frac{M_{\eta_{R}}^{2}}{M_{2}^{2}}\right)$
@@
 where $M_{2}\simeq M_{\eta_{R}}$ is taken into account.
-If we note that $h_{1}$ is supposed to be very small for the DM explanation as discussed in the previous part, the dominant contribution to this $\varepsilon$ is found to be caused by $N_{3}$ and it can be estimated by using eqs.
-(7) and (12) asdddIt should be noted that the existence of both a zero mass eigenvalue and Majorana phases of the right-handed neutrino masses are required to guarantee a non-zero $\varepsilon$ when neutrino Yukawa couplings $h_{\alpha k}$ are real [14].
+Since $h_{1}$ is supposed to be very small for the DM explanation as discussed in the previous part, the dominant contribution to this $\varepsilon$ is found to be caused by $N_{3}$ and it can be estimated by using eqs.
+(7) and (12) asfffAs addressed in the previous part, it should be noted that the existence of both a zero mass eigenvalue and Majorana phases of the right-handed neutrino masses is required to guarantee a non-zero $\varepsilon$ when neutrino Yukawa couplings $h_{\alpha k}$ are real [18].
 $\varepsilon$ $\simeq$ $\frac{h_{3}^{2}}{8\pi}\frac{{\rm Im}[(\sum_{\alpha}U_{\alpha 2}U_{\alpha 3}^{\ast})^{2}e^{i(\gamma_{2}-\gamma_{3})}]}{\sum_{\alpha}U_{\alpha 2}U_{\alpha 2}^{\ast}}F\left(\frac{M_{3}^{2}}{M_{2}^{2}},\frac{M_{\eta_{R}}^{2}}{M_{2}^{2}}\right)$ (22)
@@
 where parameters in eq.
-(14) and $\sin\theta_{ij}$ and $\alpha_{ij}$ written below it are used, and $\sin(\gamma_{2}-\gamma_{3})=1$ is assumed in the second line.
-We should note that a smaller $|\lambda_{5}|$ can make Yukawa couplings larger and then $\varepsilon$ larger but it makes the out-of-equilibrium decay of $N_{2}$ harder.
-The $N_{2}$ decay is in the equilibrium at temperature $T\geq T_{D}$. $T_{D}$ is defined by $\Gamma_{N_{2}}=H(T_{D})$ where the decay width of $N_{2}$ can be expressed as
+(14) are used and values of $\sin\theta_{ij}$ and $\alpha_{ij}$ are fixed to the ones written below eq.
+(14).
+In the second line, $\sin(\gamma_{2}-\gamma_{3})=1$ is assumed.
+If a smaller $|\lambda_{5}|$ is assumed in eq.
+(14), Yukawa couplings $h_{2,3}$ take larger values.
+In that case, $\varepsilon$ takes a larger value but the out-of-equilibrium decay of $N_{2}$ becomes harder.
+The $N_{2}$ decay at temperature $T\geq T_{D}$ is in the equilibrium if $T_{D}$ is defined by $\Gamma_{N_{2}}=H(T_{D})$. The $N_{2}$ decay width $\Gamma_{N_{2}}$ can be expressed as
 $\Gamma_{N_{2}}=\frac{q_{2}h^{2}_{2}}{16\pi}M_{2}\left(1-\frac{M_{\eta_{R}}^{2}}{M_{2}^{2}}\right)^{2}.$ (23)
-If the washout processes of the generated lepton number asymmetry are frozen out at temperature $T_{F}$, the lepton number asymmetry caused by the out-of-equilibrium decay of $N_{2}$ could escape the washout in the case $T_{D}<T_{F}$ since the dominant part of lepton number asymmetry is expected to be generated at $T~{}{{}^{<}_{\sim}}~{}T_{D}$.
-Such situation could happen if the $N_{2}$ decay is forced to be delayed and be kept in the out-of-equilibrium by making $T_{D}$ to a lower temperature.
-On the other hand, the out-of-equilibrium decay of $N_{2}$ is favored to start at the temperature $T~{}{{}^{>}_{\sim}}~{}M_{2}$ for the efficient generation of the lepton number asymmetry since the $N_{2}$ abundance can keep a large value as the relativistic particle there.
+If the washout processes of the generated lepton number asymmetry are frozen out at temperature $T_{F}$, the lepton number asymmetry caused by the out-of-equilibrium decay of $N_{2}$ could escape the washout in the case $T_{D}<T_{F}$.
+Such a situation could happen if the $N_{2}$ decay is forced to be delayed by making $\Gamma_{N_{2}}$ smaller.
 Although $h_{2}$ fixed by the neutrino oscillation data cannot realize such a situation for a TeV scale $M_{2}$, it could be realized through an additional suppression factor for $\Gamma_{N_{2}}$.
-It is caused the phase space suppression due to the strict degeneracy between $M_{2}$ and $M_{\eta_{R}}$ [16].
+It is caused by the phase space suppression due to the strict degeneracy between $M_{2}$ and $M_{\eta_{R}}$ [20].
 In fact, if their degeneracy is expressed as $M_{\eta_{R}}=(1-\delta_{2})M_{2}$ with $\delta_{2}\ll 1$, $\Gamma_{N_{2}}$ is found to be rewritten with an additional suppression factor $\delta_{2}$ as
-$\Gamma_{N_{2}}\simeq\frac{h_{2}^{2}\delta_{2}^{2}}{4\pi}M_{2}=2.9\times 10^{-12}\left(\frac{M_{2}}{3~{}{\rm TeV}}\right)\left(\frac{\delta_{2}}{10^{-4}}\right)^{2}~{}{\rm GeV},$ (24)
-where we use the values for $\sin\theta_{ij}$ and $\alpha_{ij}$ given below eq.
-(14).
-Since it shows that $\Gamma_{N_{2}}<H(M_{2})$ is satisfied for $\delta_{2}<10^{-4}$ at $M_{2}=3$ TeV, $N_{2}$ is expected to be out-of-equilibrium at $T~{}{{}^{>}_{\sim}}~{}M_{2}$.
+$\Gamma_{N_{2}}\simeq\frac{h_{2}^{2}\delta_{2}^{2}}{4\pi}M_{2}=2.9\times 10^{-12}\left(\frac{M_{2}}{3~{\rm TeV}}\right)\left(\frac{\delta_{2}}{10^{-4}}\right)^{2}~{\rm GeV}.$ (24)
+If we assume $\delta_{2}<10^{-4}$ at $M_{2}=3$ TeV, $\Gamma_{N_{2}}<H(T)$ is found to be satisfied for $T~{{}^{<}_{\sim}}~M_{2}$ and $N_{2}$ is expected to be out-of-equilibrium there.
 Thus, if the condition $T_{D}<T_{F}$ is fulfilled, the generated lepton number asymmetry $Y_{L}$ defined as $Y_{L}\equiv Y_{\ell}-Y_{\bar{\ell}}$ could be estimated as $Y_{L}=\varepsilon Y_{N_{2}}^{\rm eq}$ where $Y_{N_{2}}^{\rm eq}$ stands for the equilibrium value of $N_{2}$ as the relativistic particle.
@@
 (22).
-Unfortunately, since $\delta_{2}\ll 1$ also suppresses $\varepsilon$ heavily as found from eq.(21), this scenario cannot work.
+Unfortunately, since $\delta_{2}\ll 1$ also suppresses $\varepsilon$ heavily as found from eq.(21), this idea cannot work in the present case.
 If $M_{2}\simeq M_{1}$ or $M_{3}$ is satisfied, the $CP$ asymmetry $\varepsilon$ could be enhanced by the resonance due to a self-energy diagram mediated by $N_{1}$ or $N_{3}$.
@@
 (14).
-By calculating the relevant reaction rates of washout processes, we find that $T_{F}~{}{{}^{>}_{\sim}}~{}0.1M_{2}$ so that $\delta_{2}<3\times 10^{-5}$ guarantees $T_{D}<T_{F}$.
+By calculating the relevant reaction rates of washout processes, we find that the freezeout of the $\eta_{R}$ coannihilation occurs for $T_{F}~{{}^{>}_{\sim}}~0.1M_{2}$.
+Thus, $T_{D}<T_{F}$ is guaranteed if $\delta_{2}<3\times 10^{-5}$ is satisfied.
 This suggests that the $CP$ asymmetry $\varepsilon$ can take a value of $O(10^{-7})$ in the case $\delta_{3}=O(10^{-9})$, which makes successful leptogenesis possible as addressed below eq.
 (24).
-However, if the number density of $N_{2}$ does not decrease to the required $Y_{N_{1}}^{\infty}$, the $\eta_{R}$ produced through the $N_{2}$ decay contributes to the $N_{1}$ relic density.
+However, if the number density of $N_{2}$ does not decrease to the required $Y_{N_{1}}^{\infty}$ at the temperature where the $\eta_{R}$ coannihilation freezes out, the $\eta_{R}$ produced through the $N_{2}$ decay could contribute to the $N_{1}$ relic density.
 Thus, the estimation of the $N_{1}$ relic abundance in the previous section cannot naively applied to this scheme for leptogenesis.
-The effect of $\eta_{R}$ produced in the late time $N_{2}$ decay needs to be taken into account for the estimation of the $N_{1}$ relic abundance.
+The effect of $\eta_{R}$ produced in the late time $N_{2}$ decay needs to be taken into account in the estimation of the $N_{1}$ relic abundance.
 In order to examine its effect to the $N_{1}$ relic abundance, we solve numerically Boltzmann equations for $Y_{N_{2}}$, $Y_{L}$, $Y_{\eta_{R}}$ and $Y_{N_{1}}$.
 Using a dimensionless parameter $z_{2}$, they are given as
-$\frac{dY_{N_{2}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left\{\gamma^{S}(N_{2}\ell)+\gamma^{S}(N_{2}\eta)+\gamma^{D}(N_{2})\right\}\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right),$
-$\frac{dY_{L}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\varepsilon\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)+\left\{\sum_{k=2,3}\gamma^{D}(N_{k})+2(\gamma^{S}(\ell\eta)+\gamma^{S}(\ell\ell))\right\}\frac{Y_{L}}{Y_{\ell}^{\rm eq}}\right],$
-$\frac{dY_{\eta_{R}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\left\{2\gamma^{S}(\eta_{R}\eta_{R})+\gamma^{D}(\eta_{R})\right\}\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right)+\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)\right],$
-$\frac{dY_{N_{1}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})}\gamma^{D}(\eta_{R})\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right),$ (27)
-where $\gamma^{S}({N_{2}\ell})$ and $\gamma^{S}(N_{2}\eta)$ are reaction densities of the scattering $N_{3}\ell\leftrightarrow N_{2}\ell$ and $N_{3}\eta\leftrightarrow N_{2}\eta$ for $N_{2}$ production, and $\gamma^{S}(\ell\eta)$ and $\gamma^{S}(\ell\ell)$ are the ones for the lepton number violating scattering $\ell\eta\leftrightarrow\bar{\ell}\eta^{\dagger}$ and $\ell\ell\leftrightarrow\eta\eta$.
+$\frac{dY_{N_{2}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\Big\{\gamma^{S}(N_{2}\ell)+\gamma^{S}(N_{2}\eta)\Big\}\left(\frac{Y_{N_{2}}^{2}}{Y_{N_{2}}^{{\rm eq}2}}-1\right)+\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)\right],$
+$\frac{dY_{L}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[-\varepsilon\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)+\left\{\sum_{k=2,3}\gamma^{D}(N_{k})+2(\gamma^{S}(\ell\eta^{\dagger})+\gamma^{S}(\ell\ell))\right\}\frac{Y_{L}}{Y_{\ell}^{\rm eq}}\right],$
+$\frac{dY_{\eta_{R}}}{dz_{2}}$ $=$ $-\frac{z_{2}}{H(M_{2})s}\left[\Big\{2\gamma^{S}(\eta_{R}\eta_{R})+\gamma^{D}(\eta_{R})\Big\}\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right)-\gamma^{D}(N_{2})\left(\frac{Y_{N_{2}}}{Y_{N_{2}}^{\rm eq}}-1\right)\right],$
+$\frac{dY_{N_{1}}}{dz_{2}}$ $=$ $\frac{z_{2}}{H(M_{2})s}\gamma^{D}(\eta_{R})\left(\frac{Y_{\eta_{R}}}{Y_{\eta_{R}}^{\rm eq}}-1\right),$ (27)
+where $\gamma^{S}({N_{2}\ell})$ and $\gamma^{S}(N_{2}\eta)$ are reaction densities of the scattering $N_{3}\ell\leftrightarrow N_{2}\ell$ and $N_{3}\eta\leftrightarrow N_{2}\eta$ for the $N_{2}$ production, and $\gamma^{S}(\ell\eta^{\dagger})$ and $\gamma^{S}(\ell\ell)$ are the ones for the lepton number violating scattering $\ell\eta^{\dagger}\leftrightarrow\bar{\ell}\eta$ and $\ell\ell\leftrightarrow\eta\eta$.
 A new term is introduced to the equation for $Y_{\eta_{R}}$ to take account of the contribution from the late time decay of $N_{2}$ in comparison with the original one in eq.
 (15).
-Change of the $Y_{N_{1}}$ evolution caused by imposing the production of the required lepton number asymmetry can be found from the solutions.
+Change of the $Y_{N_{1}}$ evolution caused by imposing the production of the required lepton number asymmetry can be found from these solutions.
 Their solutions for the same values of $M_{1}$ and $(\lambda_{3},\lambda_{4})$ used in Fig.
@@
 We use the values of $M_{\eta_{R}}$, $M_{2}$ and $M_{3}$ given in (14) with $\delta_{2}=3\times 10^{-5}$ and $\delta_{3}=10^{-9}$ which are adopted referring to the above discussion.
-The figure shows that the lepton number asymmetry $|Y_{L}|$ plotted by an orange line is generated to converge to a final value at $z_{2}\sim 10$, which is larger compared with a required range to explain the baryon number asymmetry shown by dotted yellow lines.
-Since the resulting lepton number asymmetry takes a value $Y_{L}\simeq\varepsilon Y_{N_{2}}^{\rm eq}$ where $Y_{N_{2}}^{\rm eq}$ is the one of the relativistic particles, the washout of the generated lepton number asymmetry is found to be sufficiently suppressed as a result of the late time decay of $N_{2}$ as expected from the above discussion.
+The figure shows that the lepton number asymmetry $|Y_{L}|$ plotted by an orange line converges to a final value at $z_{2}\sim 10$.
+It is larger compared with a required range to explain the baryon number asymmetry shown by dotted yellow lines.
+Since the resulting lepton number asymmetry takes a value $Y_{L}\simeq\varepsilon Y_{N_{2}}^{\rm eq}$ where $Y_{N_{2}}^{\rm eq}$ is the one of the relativistic particles, the washout of the generated lepton number asymmetry is found to be sufficiently suppressed as a result of the late time decay of $N_{2}$.
+It is expected in the above discussion.
 Fig.
 2  Evolution of the lepton number asymmetry $|Y_{L}|$ and the $N_{1}$ number density $Y_{N_{1}}$ in the comoving volume.
-$|Y_{L}|$ is plotted by a orange line, which is caused by the out-of-equilibrium decay of $N_{2}$.
-$Y_{N_{2}}^{\rm eq}$ should be noted to coincide with $Y_{\eta_{R}}^{\rm eq}$ since $M_{2}\simeq M_{\eta_{R}}$ is assumed.
+$|Y_{L}|$ is plotted by an orange line, which is caused by the out-of-equilibrium decay of $N_{2}$.
+$Y_{N_{2}}^{\rm eq}$ coincides with $Y_{\eta_{R}}^{\rm eq}$ since $M_{2}\simeq M_{\eta_{R}}$ is assumed.
 Yellow dashed horizontal lines represent a region of $|Y_{L}|$ to explain the required baryon number asymmetry.
@@
 1.
-In the right panel, $(\lambda_{3},\lambda_{4})$ is changed to $(2.5,-1.0)$ from the ones used in the left panel keeping the $h_{1}$ value unchanged.
-While $Y_{N_{1}}$ becomes a little bit smaller than the required value $Y_{N_{1}}^{\infty}$ in $M_{1}=150$ GeV and 1.5 GeV, $Y_{N_{1}}+Y_{\eta_{R}}=Y_{N_{1}}^{\infty}$ is realized for $M_{1}=1.5$ TeV at $z_{2}=300$.
+In the right panel, $(\lambda_{3},\lambda_{4})$ is changed to (2.5,1.0)(2.5,-1.0) from the ones used in the left panel keeping the $h_{1}$ value unchanged.
+While $Y_{N_{1}}$ becomes a little bit smaller than the required value $Y_{N_{1}}^{\infty}$ in the cases $M_{1}=150$ GeV and 1.5 GeV, $Y_{N_{1}}+Y_{\eta_{R}}=Y_{N_{1}}^{\infty}$ is realized for $M_{1}=1.5$ TeV at $z_{2}=300$.
 It means that $Y_{N_{1}}=Y_{N_{1}}^{\infty}$ is obtained finally through the $\eta_{R}$ decay.
-On the other hand, we find from the figure that $Y_{\eta_{R}}$ leaves $Y_{\eta_{R}}^{\rm eq}$ around $z_{2}=20$, which is considered to be caused due to the break of the balance between the $\eta_{R}$ production by the $N_{2}$ decay and its disappearance due to both the decay and the coannihilation.
-It shows that the late time additional $\eta_{R}$ production through the $N_{2}$ decay affects the estimation of the $N_{1}$ relic abundance.
-It forces us to change the strength of the $\eta_{R}$ decay and the coannihilation of $\eta_{R}$ in order to realize the required $N_{1}$ abundance since we impose simultaneous explanation of the DM abundance and the baryon number asymmetry on the model.eeeSince the $N_{2}$ decay is relevant to the leptogenesis, we do not change the parameters relevant to it here.
 In the figure, $Y_{\eta_{R}}$ and $Y_{N_{1}}$ are plotted by using blue and red lines, respectively.
 Different line types are used for different masses, $M_{1}$=1.5 TeV (solid), 150 GeV (dashed), and 1.5 GeV (dotted).
-$Y_{N_{1}}^{\infty}$ which can reproduce $\Omega_{N_{1}}h^{2}=0.12$ as the present DM abundance is shown by horizontal black solid, dashed and dotted lines for each $M_{1}$.
+A $Y_{N_{1}}^{\infty}$ value which can reproduce $\Omega_{N_{1}}h^{2}=0.12$ as the present DM abundance is shown by horizontal black solid, dashed and dotted lines for each $M_{1}$.
+We find from the figure that $Y_{\eta_{R}}$ leaves $Y_{\eta_{R}}^{\rm eq}$ around $z_{2}=20$ in these cases.
+It is considered to be caused due to the break of the balance between the $\eta_{R}$ production by the $N_{2}$ decay and its disappearance due to both the decay and the coannihilation.
+It shows that the late time additional $\eta_{R}$ production through the $N_{2}$ decay affects the estimation of the $N_{1}$ relic abundance.
+It forces us to change the strength of the $\eta_{R}$ decay and the $\eta_{R}$ coannihilation from the ones in Fig.
+1 in order to realize the required $N_{1}$ relic abundancegggSince the $N_{2}$ decay is relevant to the leptogenesis, we do not change the parameters relevant to it here.
+since we impose on the model the simultaneous explanation of the DM abundance and the baryon number asymmetry.
 Since $h_{1}$ does not affect the values of $h_{2,3}$ determined by the neutrino oscillation data as long as $h_{1}$ is small enough, it can be tuned to result in the required $N_{1}$ relic abundance $Y_{N_{1}}^{\infty}$.
-By tuning $h_{1}$ suitably to $h_{1}=2.3\times 10^{-9}$ and $2.3\times 10^{-8}$ for each case $M_{1}=$150 GeV and 1.5 GeV, $Y_{N_{1}}$ is found to converge to the required $Y_{N_{1}}^{\infty}$ at $z_{2}$ where $Y_{N_{1}}=Y_{N_{2}}$ is realized.
+By tuning $h_{1}$ suitably to $h_{1}=2.3\times 10^{-9}$ and $2.3\times 10^{-8}$ for the case $M_{1}=$150 GeV and 1.5 GeV respectively, $Y_{N_{1}}$ is found to converge to the required $Y_{N_{1}}^{\infty}$ at $z_{2}$ where $Y_{N_{1}}=Y_{N_{2}}$ is realized.
 These values of $h_{1}$ are much smaller than the ones used in Fig.
 1.
-It is considered to be caused to cancel the effect of $\eta_{R}$ additionally produced through the $N_{2}$ decay by making the relative strength of the coannihilation of $\eta_{R}$ to the $\eta_{R}$ decay larger.
-The similar feature can be observed in the evolution of the number density of the keV scale $N_{1}$ also.
-For example, the DM relic abundance requires $h_{1}=6.5\times 10^{-5}$ for $M_{1}=5$ keV, and these parameters are confirmed not to affect the result of leptogenesis substantially.
-On the other hand, we find that the relation $Y_{N_{1}}+Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ cannot be changed by tuning the value of $h_{1}$ in the case $M_{1}=1.5$ TeV.
-This may be understood if we note that $Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ at $z_{2}=300$ where the coannihilation of $\eta_{R}$ is considered to be frozen out.
-It suggests that the freezeout of the coannihilation of $\eta_{R}$ plays an important role and then $\lambda_{3}$ and $\lambda_{4}$ have to be tuned especially when $M_{1}$ takes a similar order value to $M_{\eta_{R}}$.
-In the right panel, $Y_{N_{1}}$ and $Y_{\eta_{R}}$ are plotted for the case that $\lambda_{3}$ and $\lambda_{4}$ are shifted from the ones in the left panel to $(\lambda_{3},\lambda_{4})=(2.5,-1.0)$ keeping other parameters unchanged.
+It is considered to be caused to reduce the effect of $\eta_{R}$ which is additionally produced through the $N_{2}$ decay by making the relative strength of the $\eta_{R}$ coannihilation to the $\eta_{R}$ decay larger.
+The similar feature can be observed in the evolution of the number density of $N_{1}$ with the keV scale mass also.
+For example, the DM relic abundance requires $h_{1}=6.5\times 10^{-5}$ for $M_{1}=5$ keV, which can be confirmed not to affect the result of leptogenesis.
+In the case $M_{1}=1.5$ TeV, we find that the relation $Y_{N_{1}}+Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ cannot be changed by tuning the value of $h_{1}$.
+This can be understood if we note $Y_{\eta_{R}}>Y_{N_{1}}^{\infty}$ at $z_{2}=300$ where the $\eta_{R}$ coannihilation is frozen out.
+It suggests that the freezeout of the $\eta_{R}$ coannihilation plays an important role and then $\lambda_{3}$ and $\lambda_{4}$ have to be tuned when $M_{1}$ takes a similar order value to $M_{\eta_{R}}$.
+In the right panel, $Y_{N_{1}}$ and $Y_{\eta_{R}}$ are plotted for the case where $\lambda_{3}$ and $\lambda_{4}$ are shifted from the ones in the left panel to $(\lambda_{3},\lambda_{4})=(2.5,-1.0)$ keeping other parameters unchanged.
 While $Y_{N_{1}}$ in the case of $M_{1}=150$ GeV and 1.5 GeV is found to take a little bit smaller value than the one in the left panel, $Y_{N_{1}}+Y_{\eta_{R}}$ converges to $Y_{N_{1}}^{\infty}$ at $z_{2}=300$ in the case $M_{1}=1.5$ TeV.
+Since the $\eta_{R}$ coannihilation becomes stronger due to a larger value of $\lambda_{3}$, the contribution from $\eta_{R}$ is considered to be reduced.
 These examples show that $N_{1}$ in a wide mass range can explain the required relic abundance of DM in a consistent way with successful leptogenesis by tuning $h_{1}$ and $\lambda_{3,4}$ suitably.
-Couplings $\lambda_{3}$ and $\lambda_{4}$ play a crucial role especially when $M_{1}$ and $M_{\eta_{R}}$ take same order values since $Y_{N_{1}}^{\infty}~{}{{}^{>}_{\sim}}~{}Y_{\eta_{R}}^{\infty}$ should be satisfied in that case.
-We find that the lightest right-handed neutrino $N_{1}$ can be a good DM candidate consistently with leptogenesis and the model presents an interesting possibility such that three unsolved problems in the SM can be explained by low scale right-handed neutrinos simultaneously.
+Couplings $\lambda_{3}$ and $\lambda_{4}$ play a crucial role especially when $M_{1}$ and $M_{\eta_{R}}$ take the same order values since $Y_{N_{1}}^{\infty}~{{}^{>}_{\sim}}~Y_{\eta_{R}}^{\infty}$ should be satisfied in that case.
+The model presents an interesting possibility such that three unsolved problems in the SM can be explained by low scale right-handed neutrinos simultaneously.
 Finally, we should address phenomenological signatures of the model.
 Extra charged and neutral scalars $\eta^{\pm}$ and $\eta_{R,I}$ might be directly examined in the future collider experiments [35].
-However, if $N_{1}$ has a very small $h_{1}$, it might be difficult to distinguish the model from the one in which DM is $\eta_{R}$ through collider signatures.
+However, since $N_{1}$ has a very small $h_{1}$, it seems to be difficult to distinguish the model from the one in which DM is $\eta_{R}$ through collider signatures.
 The existence of right-handed neutrinos and the additional inert doublet scalar at TeV scales are expected to cause a new non-negligible contribution at one-loop level to LFV processes like $\mu\rightarrow e\gamma$ and also the electric dipole moment of electron (EDME).
@@
 Branching ratio of the LFV process $\ell_{\alpha}\rightarrow\ell_{\beta}\gamma$ in the scotogenic model is expressed as [36]
-$B(\ell_{\alpha}\rightarrow\ell_{\beta}\gamma)=\frac{3\alpha}{64\pi(G_{F}M_{\eta}^{2})^{2}}\left|\sum_{k=1}^{3}U_{\alpha k}^{\ast}U_{\beta k}h_{k}^{2}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\right|^{2},$ (28)
+$B(\ell_{\alpha}\rightarrow\ell_{\beta}\gamma)\simeq\frac{3\alpha}{64\pi(G_{F}M_{\eta}^{2})^{2}}\left|\sum_{k=1}^{3}U_{\alpha k}^{\ast}U_{\beta k}h_{k}^{2}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\right|^{2},$ (28)
 and the EDME is estimated as [18]
-$d_{e}/e=\sum_{k=1}^{3}\frac{U_{ek}^{\ast}U_{ek}h_{k}^{2}m_{e}}{16\pi^{2}M_{\eta}^{2}}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\sin\gamma_{k},$ (29)
+$d_{e}/e\simeq\sum_{k=1}^{3}\frac{U_{ek}^{\ast}U_{ek}h_{k}^{2}m_{e}}{16\pi^{2}M_{\eta}^{2}}F_{2}\left(\frac{M_{k}}{M_{\eta}}\right)\sin\gamma_{k},$ (29)
 where $F_{2}(x)$ is defined as
@@
 4 Summary
-Both existences of neutrino mass and DM are unexplained problems in the SM.
-If they are solved in a common framework, it could be a promising candidate for new physics.
+The existence of both neutrino mass and DM is unexplained problem in the SM.
+If it is solved in a common framework, it could be a promising candidate for new physics.
 The scotogenic model is such an example.
-The lightest right-handed neutrino is one of the DM candidates in this model.
-Although it could have a distinguishable character such that it cannot be found through direct detection experiments, it is known to be difficult to satisfy both constraints from the relic abundance and the LFV bounds simultaneously since they require incompatible magnitude for its Yukawa couplings.
-In order to escape this problem, additional interaction like $U(1)^{\prime}$ has been introduced to the right-handed neutrinos usually.
-Since it makes them interact with nuclei and then the direct detection experiments are relevant, the nature distinct from other ordinary WIMP candidates seems to be lost.
-In the present situation where the direct detection experiments exclude a lot of DM candidates, it may be an interesting possibility that DM is a right-handed neutrino which has no other interaction than the neutrino Yukawa couplings.
-In this study, we propose a scenario in the original scotogenic model which allows the lightest right-handed neutrino can be DM without extra interaction.
-We show that the DM abundance and the baryon number asymmetry can be simultaneously explained if we assume certain mass degeneracy among new ingredients of the model.
-Although strict mass degeneracy is required, it is noticeable that three unsolved problems in the SM can be explained through the right-handed neutrinos simultaneously.
+In this study, we propose a scenario in the original scotogenic model which allows the lightest right-handed neutrino to be DM without introducing extra interactions.
+Since this DM cannot be detected through direct detection experiments, such a model may be favorable from the present experimental status.
+We discuss remarkable features of the active neutrino mass matrix in the model and show that the right-handed neutrino with mass in a wide rage from TeV to keV can explain the required DM abundance.
+The model can also explain the baryon number asymmetry through leptogenesis based on the decay of the right-handed neutrino if we assume mass degeneracy among $Z_{2}$ odd ingredients.
+Although the required mass degeneracy is strict, it is noticeable that three unsolved problems in the SM can be explained through the right-handed neutrinos simultaneously.
 An interesting point of this scenario is that the relic abundance of DM is determined not only by the interaction of DM itself but also by the interaction of its mother particle $\eta_{R}$.
-The model gives an example that the interaction searched by the direct detection experiments of DM is irrelevant to the interaction which determines the DM abundance.
+The model gives an example that the interaction searched by the direct detection experiments of DM is irrelevant to the interaction which determines the DM relic abundance.